### v1

# Codes Correcting Transpositions of Consecutive Symbols

Abstract
The problem of correcting transpositions (or swaps) of consecutive symbols inqq-ary strings is studied. A family of codes correcting a transposition at an arbitrary location is described and proved to have asymptotically optimal redundancy. Additionally, an improved construction is given over a binary alphabet. Bounds on the cardinality of codes correcting $t=\textrm{const}$ transpositions are obtained. A lower bound on the achievable asymptotic rate of optimal codes correcting $t=\tau n$ transpositions is derived. Finally, a construction of codes correcting all possible patterns of transpositions is presented, and the corresponding lower bound on the zero-error capacity of the $q$-ary transposition channel is stated.
Keywords:
error correction bounds on codes zero-error adjacent transposition spelling error timing error bit shift permutation channel multivariate analytic combinatorics
MSC 2020: 94B25, 94B50, 94B65, 68P30, 68R05
1 Introduction and preliminaries
Communication channels with various types of reordering and synchronization errors have received increased attention from the research community lately, for the most part due to their applications in modeling modern and somewhat unconventional information transmission and storage systems. In this paper we focus on a model where the transmitted $q$-ary strings are potentially altered by transpositions of consecutive symbols. We first introduce a distance function that is appropriate for this model in the sense that it gives a characterization of the corresponding error correction problem (Section 1.1). We then present two families of codes correcting a single transposition – one over an arbitrary and another over a binary alphabet – and show that they have asymptotically optimal redundancies in the regime of growing block-length (Section 2). We further derive bounds on the cardinality of optimal codes correcting an arbitrary (but fixed) number of transpositions, and conjecture that the upper bound is asymptotically achievable up to a multiplicative constant (Section 3). In the linear regime $t=\tau n$, we derive a Gilbert–Varshamov-like lower bound on the rate of optimal codes correcting a given fraction $\tau$ of transpositions (Section 4). Since the code space under study is nonuniform, unlike in most classical settings, the derivation of such lower bounds is nontrivial and requires estimating the average ball cardinality. The methods of multivariate analytic combinatorics are employed for this purpose. Finally, we present a construction of codes correcting all possible patterns of transpositions, and state the corresponding lower bound on the zero-error capacity of the transposition channel that improves the best known bound for $q\in\{3,4\}$ (Section 5).
1.1 Model description and basic definitions
Throughout the paper, $q$ is assumed to be an integer greater than or equal to 22, and the channel alphabet is denoted by $\mathbb{Z}_{q}=\{0,1,\ldots,q-1\}$. The possible inputs and outputs from the channel are the strings from $\mathbb{Z}_{q}^{n}$, where $n\geqslant 1$ is referred to as the block-length.
For a transmitted string $\boldsymbol{x}=(x_{1},\ldots,x_{n})$, we say that a transposition [^1] Transpositions of consecutive symbols are usually called adjacent transpositions in the literature. We refer to them simply as transpositions. No confusion should arise as this is the only kind of transpositions discussed in the paper. at location $k\in\{1,\ldots,n-1\}$ has occurred in the channel if the string obtained at the channel output is $\boldsymbol{y}=(y_{1},\ldots,y_{n})$, where $y_{k}=x_{k+1}$, $y_{k+1}=x_{k}$, and $y_{i}=x_{i}$ for all $i\notin\{k,k+1\}$, i.e., if the symbols $x_{k}$ and $x_{k+1}$ have been swapped. If the transmitted string is altered by $t\geqslant 1$ transpositions, these transpositions are assumed to be applied simultaneously to disjoint pairs of symbols. In other words, if the transpositions occur at locations $k_{1},\ldots,k_{t}$, then it is understood that $k_{j+1}\geqslant k_{j}+2$ for $j=1,\ldots,t-1$. Consequently, every symbol of $\boldsymbol{x}$ can be displaced by at most one position in the channel. Here is an example of a transmitted string $\boldsymbol{x}\in\mathbb{Z}_{4}^{10}$ and the received string $\boldsymbol{y}$ obtained after $t=3$ transpositions (the pairs of symbols that are being swapped are underlined):
[…]
$\hookrightarrow\ 1\ 0\$$1\ 0\ 3\ 0\ 2\ 2\ 1\ 2\ =\ \boldsymbol{y}.$ (1b)
We denote by $B(\boldsymbol{x};t)$ the set of all strings that can be obtained by applying $t$ or fewer transpositions, of the kind just described, to the string $\boldsymbol{x}$.
A nonempty subset of $\mathbb{Z}_{q}^{n}$ is called a code of length $n$, and its elements are called codewords. A code $\mathcal{C}$ is said to be able to correct $t$ transpositions if the following condition holds:
$\forall\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}\qquad\boldsymbol{x}\neq\boldsymbol{x}^{\prime}\;\Rightarrow\;B(\boldsymbol{x};t)\cap B(\boldsymbol{x}^{\prime};t)=\emptyset.$ (2)
In words, any codeword $\boldsymbol{x}\in\mathcal{C}$, after being altered by up to $t$ transpositions in the channel, can be uniquely recovered from the received string (which belongs to $B(\boldsymbol{x};t)$) because no other codeword could have produced that string. Let $M^{\star}_{q}(n;t)$ denote the maximum cardinality of a code in $\mathbb{Z}_{q}^{n}$ correcting $t$ transpositions,
$M^{\star}_{q}(n;t)\coloneqq\max_{\mathcal{C}\subseteq\mathbb{Z}_{q}^{n},\,\mathcal{C}\,\textnormal{corrects $t$ transp.}}|\mathcal{C}|.$ (3)
We define the distance between $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{q}^{n}$ by
$d(\boldsymbol{x},\boldsymbol{y})\coloneqq\min_{t,s\colon B(\boldsymbol{x};t)\cap B(\boldsymbol{y};s)\neq\emptyset}t+s,$ (4)
with the convention $\min\emptyset=\infty$. In words, we look at all the ways a common string $\boldsymbol{z}$ can be produced by applying transpositions to both $\boldsymbol{x}$ and $\boldsymbol{y}$, and among them we choose one requiring the smallest possible cumulative number of transpositions. That number is denoted by $d(\boldsymbol{x},\boldsymbol{y})$. If no common string can be produced by applying transpositions to $\boldsymbol{x}$ and $\boldsymbol{y}$, i.e., if $B(\boldsymbol{x};t)\cap B(\boldsymbol{y};s)=\emptyset$ for all $t,s$, we set $d(\boldsymbol{x},\boldsymbol{y})=\infty$.
The ball of radius $t$ around $\boldsymbol{x}$ with respect to the distance $d(\cdot,\cdot)$ is denoted by
$\bar{B}(\boldsymbol{x};t)\coloneqq\{\boldsymbol{y}\colon d(\boldsymbol{x},\boldsymbol{y})\leqslant t\}.$ (5)
We note that $B(\boldsymbol{x};t)\subseteq\bar{B}(\boldsymbol{x};t)$, and that this containment relation is in most cases strict.
Lemma 1
For every $q\geqslant 2$ and $n\geqslant 1$, the maximum distance between two strings $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{q}^{n}$ such that $d(\boldsymbol{x},\boldsymbol{y})<\infty$ is $\max_{\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{q}^{n},\,d(\boldsymbol{x},\boldsymbol{y})<\infty}d(\boldsymbol{x},\boldsymbol{y})=n-1$.
Proof
Suppose a string $\boldsymbol{z}$ can be obtained by applying transpositions at locations $k_{1},\ldots,k_{t}$ to the string $\boldsymbol{x}$, and also by applying transpositions at locations $\ell_{1},\ldots,\ell_{s}$ to the string $\boldsymbol{y}$. If $k_{i}=\ell_{j}$, for some $i\in\{1,\ldots,t\}$, $j\in\{1,\ldots,s\}$, then $(z_{k_{i}},z_{k_{i}+1})=(x_{k_{i}+1},x_{k_{i}})=(y_{k_{i}+1},y_{k_{i}})$, and hence a common string $\boldsymbol{z}^{\prime}$ (obtained by applying a transposition at location $k_{i}$ to $\boldsymbol{z}$) can be obtained by applying t1t-1 transpositions (at locations $k_{1},\ldots,k_{i-1},k_{i+1},\ldots k_{t}$) to $\boldsymbol{x}$ and s1s-1 transpositions (at locations $\ell_{1},\ldots,\ell_{j-1},\ell_{j+1},\ldots\ell_{s}$) to $\boldsymbol{y}$. Therefore, when computing $d(\boldsymbol{x},\boldsymbol{y})$, we may, without loss of generality, assume that the sets of transposition locations $\{k_{1},\ldots,k_{t}\}$ and $\{\ell_{1},\ldots,\ell_{s}\}$ are disjoint. Since $k_{t},\ell_{s}\leqslant n-1$, this implies that $t+s\leqslant n-1$, and hence that $d(\boldsymbol{x},\boldsymbol{y})\leqslant n-1$. That this upper bound is attained is shown by the following examples. For even $n$, take $\boldsymbol{x}$ to be the alternating string $(1,0,1,0,\ldots)$ of length $n$, and $\boldsymbol{y}$ the concatenation of $(0)$, an alternating string $(0,1,0,1,\ldots)$ of length n2n-2, and $(1)$. E.g., for $n=6$, $\boldsymbol{x}=(1,0,1,0,1,0)$ and $\boldsymbol{y}=(0,0,1,0,1,1)$. Then the only common string that $\boldsymbol{x}$ and $\boldsymbol{y}$ can produce is the alternating string $\boldsymbol{z}=(0,1,0,1,\ldots)$, obtained by applying $n/2$ transpositions at locations $1,3,\ldots,n-1$ to $\boldsymbol{x}$, and n/21n/2-1 transpositions at locations $2,4,\ldots,n-2$ to $\boldsymbol{y}$. Similarly, for odd $n$, take $\boldsymbol{x}$ to be the concatenation of the alternating string $(1,0,1,0,\ldots)$ of length n1n-1 and $(0)$, and $\boldsymbol{y}$ the concatenation of $(0)$ and the alternating string $(0,1,0,1,\ldots)$. E.g., for $n=5$, $\boldsymbol{x}=(1,0,1,0,0)$ and $\boldsymbol{y}=(0,0,1,0,1)$. $\blacksquare$
The function $d(\boldsymbol{x},\boldsymbol{y})$ is symmetric, non-negative, and equals zero if and only if $\boldsymbol{x}=\boldsymbol{y}$. However, it does not satisfy the triangle inequality and is therefore not a metric. E.g., for $\boldsymbol{x}=(1,0,0,0)$, $\boldsymbol{y}=(0,0,0,1)$, $\boldsymbol{z}=(0,0,1,0)$, we have $d(\boldsymbol{x},\boldsymbol{y})=\infty$, while $d(\boldsymbol{x},\boldsymbol{z})+d(\boldsymbol{z},\boldsymbol{y})=2+1=3$. Nonetheless, it is a convenient way of measuring distance between strings in the context of the problem currently being studied and gives a sufficient condition for the error-correction capability of codes.
Define the minimum distance of a code $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ in the usual way:
$d_{\min}(\mathcal{C})\coloneqq\min_{\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C},\;\boldsymbol{x}\neq\boldsymbol{x}^{\prime}}d(\boldsymbol{x},\boldsymbol{x}^{\prime}).$ (6)
Proposition 1
Let $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$, $\mathcal{C}\neq\emptyset$. If $d_{\min}(\mathcal{C})>2t$, then $\mathcal{C}$ corrects $t$ transpositions.
Proof
By (2), the code $\mathcal{C}$ not being able to correct $t$ transpositions means that there exist distinct codewords $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}$ such that $B(\boldsymbol{x};t)\cap B(\boldsymbol{x}^{\prime};t)\neq\emptyset$. By (4) and (6), this implies that $d_{\min}(\mathcal{C})\leqslant d(\boldsymbol{x},\boldsymbol{x}^{\prime})\leqslant 2t$. $\blacksquare$
The opposite statement is in general not true. For example, let $\mathcal{C}=\{\boldsymbol{x},\boldsymbol{y}\}\subseteq\mathbb{Z}_{2}^{10}$, where $\boldsymbol{x}=(1,0,1,0,0,0,1,0,1,0)$, $\boldsymbol{y}=(0,0,1,1,0,0,0,0,1,1)$. It is easy to see that the only common string that can be produced by applying transpositions to $\boldsymbol{x}$ and $\boldsymbol{y}$ is $\boldsymbol{z}=(0,1,0,1,0,0,0,1,0,1)$, and that the only way this can be done is by applying transpositions at locations $1,3,7,9$ to $\boldsymbol{x}$, and transpositions at locations $2,8$ to $\boldsymbol{y}$. This implies that $d(\boldsymbol{x},\boldsymbol{y})=4+2=6$, and also that $B(\boldsymbol{x};3)\cap B(\boldsymbol{y};3)=\emptyset$. In other words, the code $\mathcal{C}$ corrects $t=3$ transpositions, and yet $d_{\min}(\mathcal{C})\leqslant 2t$.
Remark 1 (Transposition distance)
Another possible definition of “distance” between $\boldsymbol{x}$ and $\boldsymbol{y}$ that may seem natural in this context is the minimum number of transpositions needed to transform $\boldsymbol{x}$ into $\boldsymbol{y}$ (or, equivalently, $\boldsymbol{y}$ into $\boldsymbol{x}$), namely $d^{\prime}(\boldsymbol{x},\boldsymbol{y})\coloneqq\min\{t\colon\boldsymbol{y}\in B(\boldsymbol{x};t)\}$. However, this version of distance is not appropriate for the problem at hand because the corresponding analog of Proposition 1 does not hold. E.g., for $\boldsymbol{x}=(1,0,0)$, $\boldsymbol{y}=(0,0,1)$, we have $d^{\prime}(\boldsymbol{x},\boldsymbol{y})=\infty>2$, but $\{\boldsymbol{x},\boldsymbol{y}\}$ is not a code correcting one transposition because $\boldsymbol{z}=(0,1,0)$ can be obtained from both $\boldsymbol{x}$ and $\boldsymbol{y}$ through one transposition, i.e., $\boldsymbol{z}\in B(\boldsymbol{x};1)\cap B(\boldsymbol{y};1)$. $\blacklozenge$
1.2 A related model – successively performed transpositions
We briefly discuss here another model that is closely related to the one introduced in the previous subsection, and for which some of the results of this paper hold by an easy extension.
Namely, if the transmitted string is distorted by $t$ transpositions in the channel, one can imagine a scenario wherein these transpositions are performed in a successive manner, one after another, on pairs of symbols that are not necessarily disjoint. Here is an example with $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{4}^{10}$ and $t=3$:
[…]
In this model, pairs of symbols that are being swapped in two different steps may overlap, and thus one particular symbol of the transmitted string may end up up to $t$ positions away from its original position (see (7a)–(7c)).
Let $B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$ denote the set of all strings that can be obtained by applying $t$ or fewer successive transpositions to the string $\boldsymbol{x}$.
The definition of distance appropriate for this model is the following:
$d^{\textnormal{{\tiny(s)}}}(\boldsymbol{x},\boldsymbol{y})\coloneqq\min\!\big{\{}t\colon\boldsymbol{y}\in B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)\big{\}},$ (8)
with the convention $\min\emptyset=\infty$. In words, $d^{\textnormal{{\tiny(s)}}}(\boldsymbol{x},\boldsymbol{y})$ is the smallest number of transpositions, performed in a successive manner, that can transform $\boldsymbol{x}$ into $\boldsymbol{y}$ (or, equivalently, $\boldsymbol{y}$ into $\boldsymbol{x}$). If $\boldsymbol{y}$ cannot be obtained from $\boldsymbol{x}$ by using any number of transpositions, we set $d(\boldsymbol{x},\boldsymbol{y})=\infty$. Unlike $d(\cdot,\cdot)$, $d^{\textnormal{{\tiny(s)}}}(\cdot,\cdot)$ is a metric [^2] This metric was analyzed on the space of permutations over $\{1,\ldots,n\}$ in, e.g., [15], but can be defined for arbitrary strings.. The set $B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$ defined above is a ball of radius $t$ around $\boldsymbol{x}$ with respect to this metric. Defining the minimum distance of a code $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ with respect to $d^{\textnormal{{\tiny(s)}}}$ in the usual way (see (6)), we have the following claim.
Proposition 2
A code $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ corrects $t$ transpositions (performed in a successive manner) if and only if $d^{\textnormal{{\tiny(s)}}}_{\min}(\mathcal{C})>2t$. $\blacksquare$
The model with successive transpositions is stronger than the one with simultaneous, disjoint transpositions, in the sense that all error patterns that can be realized in the latter can also be realized in the former, i.e., $B(\boldsymbol{x};t)\subseteq B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$. (In fact, the stronger statement $\bar{B}(\boldsymbol{x};t)\subseteq B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$ is also true.) Consequently: 1.) any code construction valid for the former model is also valid for the latter, 2.) any lower bound on the cardinality of optimal codes in the former model is also valid in the latter, and 3.) any upper bound on the cardinality of optimal codes in the latter model is also valid in the former. We note that our constructions presented in Section 2 are trivially valid in both models because they concern a single transposition only, in which case there is no difference between the models. Also, the upper bound on the cardinality of optimal codes correcting $t$ disjoint transpositions presented in Theorem 3.1 is, by the above comment 3.), automatically valid for the stronger model with successive transpositions as well.
1.3 Prior work
Binary codes correcting a fixed number of successive transpositions were analyzed in [16]. Constructions of binary codes correcting a single transposition appear in some earlier works as well, e.g., in [20, 22]. We note that transpositions in binary strings are also known as bit-shifts (or peak-shifts) in the literature and are of relevance in modeling errors in magnetic recording devices. Lower bounds on achievable rates of binary codes correcting a given fraction of bit-shifts were studied in [14, 9].
Zero-error codes, i.e., codes correcting all possible patterns of transpositions, for the model studied herein (as well as for the more general model where each symbol can be displaced by at most $r$ positions) were studied in [21, 2]. Binary zero-error codes for the bit-shift channel (with an additional assumption that each 11 can be shifted by at most $\ell$ positions) were studied in [26, 19]. Zero-error codes for closely related timing channels were studied in [17, 18].
Codes for models that include other types of errors (such as deletions and substitutions) in addition to transpositions were studied in [1, 5, 6, 7, 11, 12, 25, 29, 30].
We emphasize that codes for the $q$-ary transposition channel, which are the focus of the present paper, have not been thoroughly analyzed in the literature. There are only several papers dealing with related problems, e.g., [1, 2, 3, 21, 30].
1.4 Notation
$H_{q}(x)=-x\log_{q}x-(1-x)\log_{q}(1-x)+x\log_{q}(q-1)$ is the $q$-ary entropy function. As usual, we write simply $H(x)$ for $H_{2}(x)$.
The bounds derived in the sequel are given in asymptotic form, for which we will need the following notation. For two nonnegative sequences $(a_{n})$, $(b_{n})$:
  • $\bullet$
    $a_{n}\sim b_{n}$ means $\lim_{n\to\infty}\frac{a_{n}}{b_{n}}=1$,
  • $\bullet$
    $a_{n}\lesssim b_{n}$ means $\limsup_{n\to\infty}\frac{a_{n}}{b_{n}}\leqslant 1$,
  • $\bullet$
    $a_{n}=o(b_{n})$ means $\lim_{n\to\infty}\frac{a_{n}}{b_{n}}=0$,
  • $\bullet$
    $a_{n}={\mathcal{O}}(b_{n})$ means $\limsup_{n\to\infty}\frac{a_{n}}{b_{n}}<\infty$,
  • $\bullet$
    $a_{n}=\Omega(b_{n})$ means $\liminf_{n\to\infty}\frac{a_{n}}{b_{n}}>0$, or equivalently $b_{n}={\mathcal{O}}(a_{n})$,
  • $\bullet$
    $a_{n}=\Theta(b_{n})$ means $0<\liminf_{n\to\infty}\frac{a_{n}}{b_{n}}\leqslant\limsup_{n\to\infty}\frac{a_{n}}{b_{n}}<\infty$.
2 Codes correcting a single transposition
In this section we consider the problem of correcting $t=1$ transposition.
2.1 Construction over a general alphabet
For arbitrary integers $s_{1},s_{2}$ and $p>0$, define
$\mathcal{C}_{q}(n)\coloneqq\left\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\;\colon\sum_{i=1}^{n}ix_{i}=s_{1}\pmod{2q-1},\;\;\sum_{i=1}^{n}i^{2}x_{i}=s_{2}\pmod{p}\right\}$. (9)
The dependence of the set $\mathcal{C}_{q}(n)\subseteq\mathbb{Z}_{q}^{n}$ on the parameters $s_{1},s_{2},p$ is not made explicit for reasons of notational simplicity.
Theorem 2.1
Fix integers $n\geqslant 1$, $q\geqslant 2$, $s_{1}$, $s_{2}$, and an odd prime $p\geqslant\max\{q,n\}$, such that $\mathcal{C}_{q}(n)$ is nonempty. Then the code $\mathcal{C}_{q}(n)$ corrects one transposition.
Proof
Consider a codeword $\boldsymbol{x}\in\mathcal{C}_{q}(n)$. Let $\boldsymbol{y}$ be the string obtained after a pair of consecutive symbols in $\boldsymbol{x}$ have been swapped, meaning that $(y_{k},y_{k+1})=(x_{k+1},x_{k})$ for some $k\in\{1,\ldots,n-1\}$, and $y_{i}=x_{i}$ for all $i\notin\{k,k+1\}$. The index $k$ is the transposition location that needs to be determined in order to correct the error. The receiver computes
$c_{1}\coloneqq s_{1}$ $-\sum_{i=1}^{n}iy_{i}$ $=\sum_{i=1}^{n}ix_{i}$ $-\sum_{i=1}^{n}iy_{i}$ $=x_{k+1}-x_{k}\pmod{2q-1}$, (10a)
$c_{2}\coloneqq s_{2}$ $-\sum_{i=1}^{n}i^{2}y_{i}$ $=\sum_{i=1}^{n}i^{2}x_{i}$ $-\sum_{i=1}^{n}i^{2}y_{i}$ $=(2k+1)(x_{k+1}-x_{k})\pmod{p}$. (10b)
By computing $c_{1}$ in (10a), the receiver can uniquely determine the integer $x_{k+1}-x_{k}\in\{0,\pm 1,\ldots,\pm(q-1)\}$. If $c_{1}=0$, then $\boldsymbol{y}=\boldsymbol{x}$ and there is nothing to correct. Otherwise, the receiver can uniquely determine the error location $k$ from (10b), namely
$k=2^{-1}\big{(}c_{2}(x_{k+1}-x_{k})^{-1}-1\big{)}\pmod{p}$, (11)
and recover $\boldsymbol{x}$ by swapping the symbols $y_{k}$ and $y_{k+1}$. $\blacksquare$
Instead of a prime $p$ in (9), one could take any number $n^{\prime}\geqslant n$ that is mutually prime with all numbers in the range $2,\ldots,\max\{2,q-1\}$ (so that 22 and $x_{k+1}-x_{k}$ are guaranteed to have multiplicative inverses modulo $n^{\prime}$, see (11)). We have chosen to use a prime $p$ as it makes the definition a bit simpler while not affecting the cardinality of the resulting code in the asymptotic sense (see Theorem 2.3 ahead).
2.2 Construction over a binary alphabet
In this subsection we consider the case $q=2$. Although several constructions of single-transposition-correcting binary codes were reported earlier [22, 20, 16], we present here another construction that has the best known asymptotic scaling, equaling that obtained in [16], but that is somewhat more explicit than the one from [16].
Let $\mathcal{B}(m)\subseteq\mathbb{Z}_{2}^{m}$ be a binary code of length $m$ correcting one substitution error, i.e., having minimum Hamming distance at least 33. For a string $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$, let $\boldsymbol{x}_{\textnormal{even}}=(x_{2},x_{4},\ldots)$ be the string of its even-indexed symbols. Assume for simplicity that $n$ is even, and define
$\mathcal{C}^{\prime}(n)\coloneqq\left\{\boldsymbol{x}\in\mathbb{Z}_{2}^{n}\;\colon\boldsymbol{x}_{\textnormal{even}}\in\mathcal{B}(n/2),\;\;\sum_{i=1}^{n/2}i(x_{2i}-x_{2i-1})=s\pmod{2}\right\}$, (12)
where $s\in\{0,1\}$.
Theorem 2.2
Fix integers $n\geqslant 1$ and $s$ such that $\mathcal{C}^{\prime}(n)$ is nonempty. Then the code $\mathcal{C}^{\prime}(n)$ corrects one transposition.
Proof
Consider a codeword $\boldsymbol{x}\in\mathcal{C}^{\prime}(n)$, and let $\boldsymbol{y}$ be the string obtained after a transposition at location $k$ has been applied to $\boldsymbol{x}$ so that $(y_{k},y_{k+1})=(x_{k+1},x_{k})$ for some $k\in\{1,\ldots,n-1\}$, and $y_{i}=x_{i}$ for all $i\notin\{k,k+1\}$. The receiver, given $\mathcal{C}^{\prime}(n)$ and $\boldsymbol{y}$, needs to determine $k$. The string $\boldsymbol{y}_{\textnormal{even}}=({y}_{2},y_{4},\ldots,{y}_{n})$ is the result of applying (at most) one substitution error, i.e., bit-flip, to the string $\boldsymbol{x}_{\textnormal{even}}$. Recall that the latter is a codeword of a single-substitution-correcting code $\mathcal{B}(n/2)$. Hence, if $\boldsymbol{y}_{\textnormal{even}}\in\mathcal{B}(n/2)$, there was no substitution, meaning that the transposition had no effect on the transmitted string (because $x_{k}=x_{k+1}$), and there is nothing to correct. Otherwise, if $\boldsymbol{y}_{\textnormal{even}}\notin\mathcal{B}(n/2)$, the receiver will be able to find the erroneous symbol ${y}_{2j}$. Due to the nature of transposition errors, the transposition location $k$ is either $2j$ or 2j12j-1, and hence all that is left for the receiver to determine is whether $k$ is even or odd. To that end, the receiver computes
$c\coloneqq s-\sum_{i=1}^{n/2}i(y_{2i}-y_{2i-1})=\sum_{i=1}^{n/2}i(x_{2i}-x_{2i-1})-\sum_{i=1}^{n/2}i(y_{2i}-y_{2i-1})\pmod{2}$. (13)
It can be verified directly that, if k=2j1k=2j-1, then the sum in (13) equals $(k+1)(x_{k+1}-x_{k})$, and hence, since $x_{k+1}-x_{k}=\pm 1$, $c=0$. Similarly, if $k=2j$, the sum in (13) equals $(k+1)(x_{k}-x_{k+1})$, and hence $c=1$. Therefore, the calculated bit $c$ contains the information about the parity of $k$, and the transposition location can be computed as k=2j1+ck=2j-1+c. $\blacksquare$
2.3 Bounds
Theorem 2.3 below gives asymptotic bounds on the maximum possible cardinality of a code in $\mathbb{Z}_{q}^{n}$ correcting one transposition, namely $M^{\star}_{q}(n;1)$, and implies that the codes $\mathcal{C}_{q}(n)$ described in the previous subsection are asymptotically optimal in the sense that
$M^{\star}_{q}(n;1)=\Theta\!\left(|\mathcal{C}_{q}(n)|\right)=\Theta\Big{(}\frac{q^{n}}{n}\Big{)}.$ (14)
Before establishing the theorem, we state one auxiliary fact that will be needed in the proof.
By a run in a string $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$ we mean a substring of identical symbols that is delimited on both sides either by a different symbol or by the end of the string. The number of runs in $\boldsymbol{x}$ is denoted by $\operatorname{run}(\boldsymbol{x})$. For example, the string $\boldsymbol{x}=(0,3,3,3,0,0,1,2,2)\in\mathbb{Z}_{4}^{9}$ has $\operatorname{run}(\boldsymbol{x})=5$ runs.
Lemma 2
For fixed $\rho\in[0,1]$,
$\lim_{n\to\infty}\frac{1}{n}\log_{q}\!\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\;\colon\operatorname{run}(\boldsymbol{x})\leqslant\rho n\}\right|=\begin{cases}H_{q}(\rho),&0\leqslant\rho<1-q^{-1}\\ 1,&1-q^{-1}\leqslant\rho\leqslant 1\end{cases},$ (15)
where $H_{q}(\cdot)$ is the $q$-ary entropy function.
Proof
Let $S_{q}(n,r)$ denote the number of strings from $\mathbb{Z}_{q}^{n}$ having exactly $r$ runs of identical symbols, so that
$\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\;\colon\operatorname{run}(\boldsymbol{x})\leqslant\rho n\}\right|=\sum_{r=1}^{\lfloor\rho n\rfloor}S_{q}(n,r).$ (16)
By a direct counting argument we find that
$S_{q}(n,r)=\binom{n-1}{r-1}q(q-1)^{r-1}$ (17)
and hence, by Stirling’s approximation,
$\lim_{n\to\infty}\frac{1}{n}\log_{q}S_{q}(n,\lfloor\rho n\rfloor)=H_{q}(\rho).$ (18)
Since the quantity we are interested in, expressed in (16), can be sandwiched between
$\max_{r\leqslant\rho n}S_{q}(n,r)\;\leqslant\;\sum_{r=1}^{\lfloor\rho n\rfloor}S_{q}(n,r)\;\leqslant\;\rho n\cdot\max_{r\leqslant\rho n}S_{q}(n,r),$ (19)
its exponential growth-rate (to base $q$) is
$\lim_{n\to\infty}\frac{1}{n}\log_{q}\sum_{r=1}^{\lfloor\rho n\rfloor}S_{q}(n,r)=\max_{0\leqslant\rho^{\prime}\leqslant\rho}H_{q}(\rho^{\prime}).$ (20)
That this is equal to the right-hand side of (15) follows from the fact that $H_{q}(\rho)$ is a concave function of $\rho\in[0,1]$ maximized at $\rho=1-q^{-1}$, the maximum being $H_{q}(1-q^{-1})=1$. $\blacksquare$
Lemma 2 implies that, for any fixed $\epsilon>0$, the number of strings having at most $(1-q^{-1}-\epsilon)n$ runs is exponentially smaller than the number of remaining strings, namely
$\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\;\colon\operatorname{run}(\boldsymbol{x})\leqslant(1-q^{-1}-\epsilon)n\}\right|=q^{nH_{q}(1-q^{-1}-\epsilon)+o(n)},$ (21)
where $H(1-q^{-1}-\epsilon)<1$. A very useful consequence of this fact is that one can disregard such strings in the asymptotic analysis of optimal codes. Namely, by noticing that any code correcting 22 substitutions, i.e., having minimum Hamming distance >4>\!4, is also capable of correcting one transposition (because a transposition affects two symbols in the original string), and by using the best known such codes, it follows that $M^{\star}_{q}(n;1)=\Omega(\frac{q^{n}}{n^{4}})$ and hence
$M^{\star}_{q}(n;1)=\Theta\!\left(q^{n+o(n)}\right).$ (22)
Relations (21) and (22) imply that, for an optimal single-transposition-correcting code, throwing out its codewords that have at most $(1-q^{-1}-\epsilon)n$ runs affects its cardinality only by an asymptotically negligible amount.
Theorem 2.3
For any fixed $q\geqslant 2$, as $n\to\infty$,
$\frac{q^{n}}{(2q-1)\,n}\;\lesssim\;M^{\star}_{q}(n;1)\;\lesssim\;\frac{q^{n+1}}{(q-1)\,n}.$ (23)
In the binary case ($q=2$), the lower bound can be improved to $\frac{2^{n}}{n}$.
Proof
The lower bound in (23) follows from the construction presented in Section 2.1. Namely, by the pigeon-hole principle there exist parameters $s_{1}\in\{1,\ldots,2q-1\}$ and $s_{2}\in\{1,\ldots,p\}$ such that the code $\mathcal{C}_{q}(n)$ has at least $\frac{q^{n}}{(2q-1)\,p}$ codewords (see (9)). Since the smallest prime $p$ greater than or equal to $n$ satisfies $p=n+o(n)$ when $n\to\infty$ [4, Section 1.4.1], the lower bound in (23) follows.
Similarly, the lower bound of $\frac{2^{n}}{n}$ for the binary case follows from the construction presented in Section $2.$ (We note that this bound was also reported in [16, Theorem 11].) Namely, by the pigeon-hole principle, for at least one value of $s\in\{0,1\}$ the code $\mathcal{C}^{\prime}(n)$ has at least $\frac{|\mathcal{B}(n/2)|\cdot 2^{n/2}}{2}$ codewords. The result then follows by recalling that there exist binary single-substitution-correcting codes $\mathcal{B}(m)$ (e.g., Hamming codes) satisfying $|\mathcal{B}(m)|\gtrsim\frac{2^{m}}{m}$.
The upper bound in (23) is obtained by a packing argument. Let $\mathcal{C}^{\star}_{q}(n)$ be an optimal code correcting one transposition, meaning that $|\mathcal{C}^{\star}_{q}(n)|=M^{\star}_{q}(n;1)$. Recall that $B(\boldsymbol{x};1)$ denotes the set of all strings that can be obtained by applying at most one transposition to a string $\boldsymbol{x}$, and note that
$|B(\boldsymbol{x};1)|=\operatorname{run}(\boldsymbol{x}).$ (24)
This is because only the transpositions occurring at the boundaries of consecutive runs in $\boldsymbol{x}$ will actually alter $\boldsymbol{x}$, and hence there are $\operatorname{run}(\boldsymbol{x})-1$ strings different from $\boldsymbol{x}$ that can be obtained in this way. Accounting for $\boldsymbol{x}$ itself, (24) follows. We further have
$\sum_{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n)\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n}|B(\boldsymbol{x};1)|\;\leqslant\;\sum_{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n)}|B(\boldsymbol{x};1)|\;\leqslant\;|\mathbb{Z}_{q}^{n}|\;=\;q^{n},$ (25)
where the second inequality follows from the fact that the sets $B(\boldsymbol{x};1)$ and $B(\boldsymbol{x}^{\prime};1)$ are disjoint for any two distinct codewords $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}^{\star}_{q}(n)$ (see (2)). From (24) and (25) we then obtain
$(1-q^{-1}-\epsilon)n\cdot\left|\left\{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n)\;\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n\right\}\right|\;<\;q^{n}.$ (26)
From the discussion after the proof of Lemma 2 we know that we can safely ignore the codewords with at most $(1-q^{-1}-\epsilon)n$ runs in the asymptotic analysis, meaning that
$\left|\left\{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n)\;\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n\right\}\right|\;\sim\;\left|\mathcal{C}^{\star}_{q}(n)\right|\;=\;M^{\star}_{q}(n;1).$ (27)
Now (26) and (27) imply
$\left(1-q^{-1}-\epsilon\right)n\cdot M^{\star}_{q}(n;1)\;\lesssim\;q^{n}.$ (28)
Since (28) holds for any $\epsilon>0$, the upper bound from (23) is established. $\blacksquare$
Consequently, as $n\to\infty$,
$M^{\star}_{q}(n;1)\;=\;\Theta\Big{(}\frac{q^{n}}{n}\Big{)},$ (29)
or, written differently,
$\log_{q}M^{\star}_{q}(n;1)\;=\;n-\log_{q}n+{\mathcal{O}}(1).$ (30)
Resilience to one transposition error requires $\log_{q}n+{\mathcal{O}}(1)$ redundant symbols.
3 Codes correcting $t=\textnormal{const}$ transpositions
In this section we present bounds on the cardinality of optimal codes in $\mathbb{Z}_{q}^{n}$ correcting $t$ transpositions (Theorem 3.1 ahead). The upper bound generalizes to arbitrary $t$ the bound stated in Theorem 2.3 for the special case $t=1$.
Recall that $B(\boldsymbol{x};t)$ denotes the set of all strings that can be obtained by applying up to $t$ transpositions to the string $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$, and $\bar{B}(\boldsymbol{x};t)$ denotes a ball (with respect to distance $d(\cdot,\cdot)$) of radius $t$ around $\boldsymbol{x}$. We first state bounds on these quantities that will be useful in what follows.
Lemma 3
For every $q\geqslant 2$, $n\geqslant 1$, $t\geqslant 1$, and $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$,
$\sum_{u=0}^{t}\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor}{u}\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor-2u-1}{t-u}\;\leqslant\;|B(\boldsymbol{x};t)|\;\leqslant\;\sum_{s=0}^{t}\binom{n-s}{s}$. (31)
Proof
To derive a lower bound on the cardinality of $B(\boldsymbol{x};t)$, consider the odd-numbered runs (first, third, etc.) in $\boldsymbol{x}$. Every transposition in $\boldsymbol{x}$ occurs either at the left or at the right boundary of an odd-numbered run, and so every pattern of exactly $t$ transpositions can be described by assuming that $u$ transpositions occur at the right boundary of such a run in the first phase (i.e., we select $u$ odd-numbered runs and swap their right-most symbols with the left-most symbols of the even-numbered runs immediately to their right), after which tut-u transpositions occur at the left boundary of such a run in the second phase, for some $u=0,1,\ldots,t$. The former can happen in $\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor}{u}$ ways, and the latter in at least $\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor-2u-1}{t-u}$ ways, implying the left-hand inequality in (31). (In the second phase we exclude the first run, because it has no boundary on the left. Also, if the $i$’th run was selected for the transposition in the first phase, we exclude in the second phase both the $i$’th run, because it is possible that it is of length 11 whence its left and right boundaries coincide, and the $(i+2)$’th run, because it is possible that the $(i+1)$’th run in between them was of length 11. Hence the 2u1-2u-1 term.)
To prove the stated upper bound on the cardinality of $B(\boldsymbol{x};t)$, it suffices to show that there are at most $\binom{n-s}{s}$ strings that can be obtained by applying exactly $s$ transpositions to $\boldsymbol{x}$. Since the transposition locations $k_{1},\ldots,k_{s}$ belong to $\{1,\ldots,n-1\}$ and satisfy $k_{j+1}>k_{j}+1$, every set of $s$ transposition locations corresponds uniquely to a binary string with $s$ 11’s and nsn-s 0’s such that there is at least one 0 after every 11. The number of such strings can be found by counting the number of ways to distribute n2sn-2s 0’s in $s+1$ blocks (where the blocks are separated by pairs of symbols $1,0$), which is $\binom{(n-2s)+(s+1)-1}{(s+1)-1}=\binom{n-s}{s}$. $\blacksquare$
Lemma 4
For every $q\geqslant 2$, $n\geqslant 1$, $t\geqslant 1$, and $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$,
$|\bar{B}(\boldsymbol{x};t)|$ $\;\leqslant\;\sum_{v=0}^{2t/3}\binom{2(t-v)}{v}|B(\boldsymbol{x};t-v)|$ (32a)
$\;\leqslant\;\Big{(}\frac{2t}{3}+1\Big{)}\max_{v}\binom{2(t-v)}{v}|B(\boldsymbol{x};t-v)|.$ (32b)
Proof
We can think of $\bar{B}(\boldsymbol{x};t)$ as the set of all strings that can be obtained through a two-step process: first apply up to $u$ transpositions to $\boldsymbol{x}$, resulting in a string $\boldsymbol{y}$, and then apply $v$ transpositions to $\boldsymbol{y}$, where $u+v=t$.
Taking $u=t$ in the first step (and hence $v=0$ in the second), we get exactly $B(\boldsymbol{x};t)$, i.e., the set of strings that can be obtained from $\boldsymbol{x}$ with at most $t$ transpositions. This is the summand $v=0$ in (32a).
Consider now the case u=t1u=t-1, and let $\boldsymbol{y}\in B(\boldsymbol{x};t-1)$ be a string obtained from $\boldsymbol{x}$ in the first step. In the second step we are allowed to choose $v=1$ pair of consecutive symbols in $\boldsymbol{y}$ to swap. If this pair is disjoint from all the pairs chosen in the first step, then the resulting string $\boldsymbol{z}$ can equivalently be obtained directly from $\boldsymbol{x}$ via up to $t$ disjoint transpositions. In other words, $\boldsymbol{z}\in B(\boldsymbol{x};t)$ and this string was therefore already accounted for in the first case described in the previous paragraph. Also, if the pair chosen in the second step is identical to a pair chosen in the first step, then the transposition from the second step will annul the transposition from the first step and, again, $\boldsymbol{z}\in B(\boldsymbol{x};t)$, i.e., we are back to the previous case. Therefore, in the second step we can, without loss of generality, consider only pairs of symbols that partially overlap with a pair chosen in the first step, either on its left or on its right symbol. (E.g., suppose that $(a,b,c,d)$ is a substring of $\boldsymbol{x}$, and that the pair $(b,c)$ was chosen in the first step, so that the corresponding substring of $\boldsymbol{y}$ is $(a,c,b,d)$. Then the pairs in $\boldsymbol{y}$ that partially overlap with the pair chosen in the first step, one of which we can choose in the second step, are $(a,c)$ and $(b,d)$.) Since at most $u$ pairs were chosen in the first step, there are at most $2u$ pairs to select one from in the second step. In conclusion, there are $|B(\boldsymbol{x};u)|$ different strings $\boldsymbol{y}$ that can be obtained in the first step, and for each such $\boldsymbol{y}$ there are at most $\binom{2u}{1}=2u$ different strings $\boldsymbol{z}$ that can be obtained in the second step.
The analysis in the general case is the same: there are $|B(\boldsymbol{x};u)|$ different strings $\boldsymbol{y}$ that can be obtained in the first step, and for each such $\boldsymbol{y}$ there are at most $\binom{2u}{v}$ different strings $\boldsymbol{z}$ that can be obtained in the second step. $\blacksquare$
Theorem 3.1
For any fixed $q\geqslant 2$ and $t\geqslant 1$, as $n\to\infty$,
$\frac{(2t)!\,q^{n}}{(4t/3+1)\,n^{2t}}\;\lesssim\;M^{\star}_{q}(n;t)\;\lesssim\;\frac{t!\,q^{n+t}}{(q-1)^{t}\,n^{t}}.$ (33)
In the binary case ($q=2$), the lower bound can be improved to $c_{t}\frac{2^{n+t}}{n^{t}}$, where $c_{1}=\frac{1}{2}$, $c_{2}=\frac{1}{3}$, and $c_{t}=\frac{1}{2t+1}$ for $t\geqslant 3$.
Proof
As noted in Section 1.1, the distance function $d(\cdot,\cdot)$ is not a metric, and hence $\bar{B}(\boldsymbol{x};t)$ is not a metric ball. Nonetheless, the same argument used to obtain the Gilbert–Varshamov bound in classical settings [8, 28] can be applied here as well. Namely, by a greedy procedure one is guaranteed to find a collection of strings $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ satisfying the following two conditions:
  • 1.)
    $d(\boldsymbol{x},\boldsymbol{x}^{\prime})>2t$ (i.e., $\boldsymbol{x}^{\prime}\notin\bar{B}(\boldsymbol{x};2t)$, see (5)) for any two distinct $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}$,
  • 2.)
    $\bigcup_{\boldsymbol{x}\in\mathcal{C}}\bar{B}(\boldsymbol{x};2t)\supseteq\mathbb{Z}_{q}^{n}$.
The first condition implies, by Proposition 1, that $\mathcal{C}$ is a code correcting $t$ transpositions. The second condition, stating that balls of radius $2t$ around the codewords cover the whole space, implies that $\sum_{\boldsymbol{x}\in\mathcal{C}}|\bar{B}(\boldsymbol{x};2t)|\geqslant q^{n}$, and hence
$|\mathcal{C}|\;\geqslant\;\frac{q^{n}}{\max_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};2t)|}.$ (34)
By Lemmas 3 and 4, we have
$\max_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};2t)|\;\leqslant\;\Big{(}\frac{4t}{3}+1\Big{)}\max_{v}\binom{2(2t-v)}{v}\sum_{s=0}^{2t-v}\binom{n-s}{s}$. (35)
Since $\binom{n-s}{s}\sim\frac{n^{s}}{s!}$ when $n\to\infty$, the dominant term on the right-hand side of (35) is obtained for $v=0$ and $s=2t$, implying
$\max_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};2t)|\;\lesssim\;\Big{(}\frac{4t}{3}+1\Big{)}\frac{n^{2t}}{(2t)!}.$ (36)
By (34) and (36), we have established the existence of a family of $t$-transposition-correcting codes of cardinality $\gtrsim\frac{(2t)!\,q^{n}}{(4t/3+1)\,n^{2t}}$.
The lower bound of $c_{t}\frac{2^{n+t}}{n^{t}}$ for the binary case was derived in [16, Theorem 11] for the stronger model with successive transpositions, and is therefore automatically valid for the model with disjoint transpositions (see Section 1.2).
In order to show the upper bound in (33), an asymptotic lower bound on the cardinality of $B(\boldsymbol{x};t)$ will be needed. Using Lemma 3, for strings satisfying $\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n$, $|B(\boldsymbol{x};t)|$ can be asymptotically lower-bounded by
$\gtrsim\;\sum_{u=0}^{t}\binom{(1-q^{-1}-\epsilon)n/2}{u}\binom{(1-q^{-1}-\epsilon)n/2}{t-u}$ (37a)
$\sim\;\sum_{u=0}^{t}\frac{(1-q^{-1}-\epsilon)^{u}\,n^{u}}{2^{u}\,u!}\frac{(1-q^{-1}-\epsilon)^{t-u}\,n^{t-u}}{2^{t-u}\,(t-u)!}$ (37b)
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{2^{t}}\,\sum_{u=0}^{t}\frac{1}{u!\;(t-u)!}$ (37c)
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{2^{t}\,t!}\,\sum_{u=0}^{t}\binom{t}{u}$ (37d)
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{t!}.$ (37e)
(In (37b) we used the fact that $\binom{n}{s}\sim\frac{n^{s}}{s!}$ when $s$ is fixed and $n\to\infty$.) The rest of the proof is similar to that of Theorem 2.3. One first notes that, by (21), disregarding all codewords having at most $(1-q^{-1}-\epsilon)n$ runs incurs no penalty in terms of the asymptotic scaling. Further, by (2), the sets $B(\boldsymbol{x};t)$ centered at codewords of an arbitrary code $\mathcal{C}$ correcting $t$ transpositions do not overlap, and hence $\sum_{\boldsymbol{x}\in\mathcal{C}}|B(\boldsymbol{x};t)|\leqslant|\mathbb{Z}_{q}^{n}|$. These facts, together with (37), imply
$M^{\star}_{q}(n;t)\cdot\frac{1}{t!}\left(1-q^{-1}-\epsilon\right)^{\!t}n^{t}\;\lesssim\;q^{n}.$ (38)
Since this relation holds for any $\epsilon>0$, the upper bound in (33) follows. $\blacksquare$
Therefore, attaining resilience to $t$ transpositions requires at least $t\log_{q}n+{\mathcal{O}}(1)$ and at most $2t\log_{q}n+{\mathcal{O}}(1)$ redundant symbols. Based on what is known in the special case $t=1$ (see (29)), as well as in the case of general $t$ and $q=2$, namely $M^{\star}_{2}(n;t)=\Theta(\frac{2^{n}}{n^{t}})$, it is reasonable to conjecture that
$M^{\star}_{q}(n;t)\;\stackrel{{\scriptstyle?}}{{=}}\;\Theta\Big{(}\frac{q^{n}}{n^{t}}\Big{)},$ (39)
i.e., that the redundancy of $t\log_{q}n+{\mathcal{O}}(1)$ is achievable for any fixed alphabet size $q$. At present, however, we do not have a lower bound on the cardinality of optimal codes that would imply this result when $q>2$ and $t>1$.
Remark 2 (Lower bound for the binary case)
The bound $M^{\star}_{2}(n;t)\gtrsim c_{t}\frac{2^{n+t}}{n^{t}}$ was derived in [16, Theorem 11]333[16, Theorem 11] states a slightly worse constant $c_{2}=\frac{1}{4}$ than Theorem 3.1 of the present paper. It can be shown based on [27, Theorem 3] that $c_{2}$ can be improved to $\frac{1}{3}$. by exhibiting a geometric interpretation of binary codes correcting successive transposition (a.k.a. bit-shift) errors as packings in the Manhattan metric, and using the densest known packings. Unfortunately, the mentioned interpretation fails when $q>2$ and thus the construction that implies this lower bound cannot be generalized to larger alphabets in any obvious way. $\blacklozenge$
4 Codes correcting $t=\tau n$ transpositions
In this section we consider the regime where the number of transpositions grows linearly with the block-length. We derive a lower bound on the asymptotic rate of optimal codes in $\mathbb{Z}_{q}^{n}$ correcting $t=\tau n$ transpositions, for any fixed $\tau\in[0,\frac{1}{2}]$ (Theorem 4.1 ahead). The bound will in fact be expressed as a function of the minimum distance, as is more common in the literature, but it can easily be written in terms of the number of correctable transpositions as well. Namely, Proposition 1 guarantees that codes of relative distance $\delta$ can correct a fraction of $\tau=\frac{\delta}{2}$ transpositions.
4.1 Generalized Gilbert–Varshamov bound
The standard Gilbert–Varshamov bound states that the cardinality of an optimal code of minimum distance $d$ is lower-bounded by the ratio of the cardinality of the code space and the cardinality of a ball of radius d1d-1 [8, 28]. The bound was later generalized [13, 10] to include nonuniform spaces in which the cardinality of a ball of given radius depends on its center. Namely, it was proven in [10] that the cardinality of an optimal code of minimum distance $d$ is lower-bounded by the ratio of the cardinality of the code space and the average cardinality of a ball of radius d1d-1.
Let $\bar{M}_{q}^{\star}(n;d)$ denote the maximum cardinality of a code in $\mathbb{Z}_{q}^{n}$ of minimum distance $d$ with respect to the distance function (4). As already noted at the beginning of the proof of Theorem 3.1, even though the distance function (4) is not a metric, the Gilbert–Varshamov bound still holds in the case analyzed in the present paper, implying that
$\bar{M}_{q}^{\star}(n;d)\;\geqslant\;\frac{q^{n}}{\frac{1}{q^{n}}\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};r)|}\;=\;\frac{q^{2n}}{\bar{T}_{q}(n;d-1)},$ (40)
where
$\bar{T}_{q}(n;r)\coloneqq\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};r)|$ (41)
is the “total ball size” of radius $r$. By Proposition 1, we have $M_{q}^{\star}(n;t)\geqslant\bar{M}_{q}^{\star}(n;2t)$, so (40) is automatically a lower bound on $M_{q}^{\star}(n;\lfloor\frac{d-1}{2}\rfloor)$.
Let $\bar{R}_{q}^{\star}(\delta)$ be the largest asymptotic rate of codes in $\mathbb{Z}_{q}^{n}$ of minimum distance $\delta n$, expressed in bits per symbol,
$\bar{R}_{q}^{\star}(\delta)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}\bar{M}_{q}^{\star}(n;\delta n),$ (42)
and denote the exponential growth rate (to base 22) of $\bar{T}_{q}(n;\rho n)$ by
$\bar{\beta}_{q}(\rho)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}\bar{T}_{q}(n;\rho n).$ (43)
Then (40) implies that
$\bar{R}^{\star}_{q}(\delta)\;\geqslant\;2\log_{2}(q)-\bar{\beta}_{q}(\delta).$ (44)
Therefore, an upper bound on the function $\bar{\beta}_{q}(\cdot)$ would imply a lower bound on the achievable asymptotic rates of transposition-correcting codes.
4.2 Estimating the average cardinality of $\bar{B}(\boldsymbol{x};r)$
Let $A(\boldsymbol{x};r)$ be the set of strings that can be obtained from $\boldsymbol{x}$ by exactly $r$ transpositions. Denote
$P_{q}(n;r)\coloneqq\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|A(\boldsymbol{x};r)|,$ (45)
and
$\alpha_{q}(\rho)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}P_{q}(n;\rho n).$ (46)
Note that the maximum number of disjoint transpositions $r$ is $\frac{n}{2}$, so the domain of $\alpha_{q}(\cdot)$ is $0,\frac{1}{2}]$.
Lemma 5
For $q\geqslant 2$ and $\rho\in[0,\frac{1}{2}]$,
$\alpha_{q}(\rho)=(1-\rho)H\Big{(}\frac{\rho}{1-\rho}\Big{)}+(1-\rho)\log_{2}(q)+\rho\log_{2}(q-1),$ (47)
where $H(\cdot)$ is the binary entropy function. The function $\alpha_{q}(\rho)$ is concave and maximized for
$\rho^{*}=\frac{1}{2}\left(1-\sqrt{\frac{q}{5q-4}}\right).$ (48)
Proof
We shall derive the generating function of the bivariate sequence $(P_{q}(n;r))_{n,r}$ and use the methods of analytic combinatorics [24] to extract the exponent $\alpha_{q}(\rho)$ from it.
Consider the two cases:
  • $\bullet$
    If the last two symbols of $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$ are different, i.e., $\boldsymbol{x}=\boldsymbol{x}^{\prime}yy^{\prime}$ with $\boldsymbol{x}^{\prime}\in\mathbb{Z}_{q}^{n-2}$, $^{\prime}\in\mathbb{Z}_{q}, yyy\neq y^{\prime}$, then
    $|A(\boldsymbol{x};r)|=|A(\boldsymbol{x}^{\prime}y;r)|+|A(\boldsymbol{x}^{\prime};r-1)|.$ (49a)
    The first term corresponds to the case when all the transpositions occur in the substring $\boldsymbol{x}^{\prime}y$, and the second term to the case when up to r1r-1 transpositions occur in the substring $\boldsymbol{x}^{\prime}$ and one transposition occurs in $yy^{\prime}$.
  • $\bullet$
    If the last two symbols of $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$ are the same, i.e., $\boldsymbol{x}=\boldsymbol{x}^{\prime}yy$, then
    $|A(\boldsymbol{x};r)|=|A(\boldsymbol{x}^{\prime}y;r)|.$ (49b)
We now have
$F_{q}(z,u)$ $\coloneqq\sum_{n,r\geqslant 0}P_{q}(n;r)z^{n}u^{r}$
$=\sum_{n,r}\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|A(\boldsymbol{x};r)|z^{n}u^{r}$
$=1+\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{y\in\mathbb{Z}_{q}}\sum_{\boldsymbol{x}^{\prime}yy\in\mathbb{Z}_{q}^{n}}|A(\boldsymbol{x}^{\prime}y;r)|z^{n}u^{r}$
$\phantom{=\ 1}+\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{y\in\mathbb{Z}_{q}}\sum_{y^{\prime}\in\mathbb{Z}_{q}\setminus\{y\}}\sum_{\boldsymbol{x}^{\prime}yy^{\prime}\in\mathbb{Z}_{q}^{n}}(|A(\boldsymbol{x}^{\prime}y;r)|+|A(\boldsymbol{x}^{\prime};r-1)|)z^{n}u^{r}$
$=1+\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{\boldsymbol{x}^{\prime}y\in\mathbb{Z}_{q}^{n-1}}|A(\boldsymbol{x}^{\prime}y;r)|z^{n}u^{r}$
$\phantom{=\ 1}+(q-1)\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{\boldsymbol{x}^{\prime}y\in\mathbb{Z}_{q}^{n-1}}|A(\boldsymbol{x}^{\prime}y;r)|z^{n}u^{r}$
$\phantom{=\ 1}+q(q-1)\sum_{n\geqslant 1}\sum_{r\geqslant 0}\sum_{\boldsymbol{x}^{\prime}\in\mathbb{Z}_{q}^{n-2}}|A(\boldsymbol{x}^{\prime};r-1)|z^{n}u^{r}$
$=1+zF_{q}(z,u)+(q-1)zF_{q}(z,u)+q(q-1)z^{2}uF_{q}(z,u)$
$=1+(qz+q(q-1)z^{2}u)F_{q}(z,u),$ (50)
and hence
$F_{q}(z,u)=\frac{1}{1-qz-q(q-1)z^{2}u}.$ (51)
Denote by $G_{q}(z,u)=1-qz-q(q-1)z^{2}u$ the denominator in (51). We now use [23, Theorem 1.3], which implies that the exponential growth rate of $P_{q}(n;\rho n)$ equals
$\alpha_{q}(\rho)=-\log_{2}(z^{*})-\rho\log_{2}(u^{*}),$ (52)
where $z^{*},u^{*}$ are the unique positive solutions to the equations $G_{q}(z,u)=0$ and $\rho z\frac{\partial G_{q}(z,u)}{\partial z}=u\frac{\partial G_{q}(z,u)}{\partial u}$, i.e.,
$1-qz-q(q-1)z^{2}u$ $=0$ (53a)
$\rho(qz+2q(q-1)z^{2}u)$ $=q(q-1)z^{2}u.$ (53b)
Solving these equations, we obtain
$z^{*}=\frac{1-2\rho}{q(1-\rho)},\quad u^{*}=\frac{q\rho(1-\rho)}{(q-1)(1-2\rho)^{2}},$ (54)
and plugging the solutions back into (52) gives
$\alpha_{q}(\rho)=$ $-\rho\log_{2}(\rho)-(1-2\rho)\log_{2}(1-2\rho)+(1-\rho)\log_{2}(1-\rho)$
$+(1-\rho)\log_{2}(q)+\rho\log_{2}(q-1)$
$=$ $\ (1-\rho)H\Big{(}\frac{\rho}{1-\rho}\Big{)}+(1-\rho)\log_{2}(q)+\rho\log_{2}(q-1).$ (55)
The second part of the statement is shown by calculating the derivatives of the function $\alpha_{q}(\rho)$. $\blacksquare$
Recall that $B(\boldsymbol{x};r)$ represents the set of strings that can be obtained from $\boldsymbol{x}$ by at most $r$ disjoint transpositions. Denote
$T_{q}(n;r)\coloneqq\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|B(\boldsymbol{x};r)|$ (56)
and
$\beta_{q}(\rho)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}T_{q}(n;\rho n).$ (57)
Lemma 6
For $q\geqslant 2$ and $\rho\geqslant 0$, we have
$\beta_{q}(\rho)=\max_{\rho^{\prime}\in[0,\rho]}\alpha_{q}(\rho^{\prime})=\begin{cases}\alpha_{q}(\rho),&\rho\leqslant\rho^{*}\\ \alpha_{q}(\rho^{*}),&\rho\geqslant\rho^{*}\end{cases}$ (58)
where $\alpha_{q}(\cdot)$ and $\rho^{*}$ are given in Lemma 5.
Proof
Since $B(\boldsymbol{x};r)=\bigcup_{r^{\prime}=0}^{r}A(\boldsymbol{x};r^{\prime})$, we have
$\max_{r^{\prime}\in\{0,\ldots,r\}}|A(\boldsymbol{x};r^{\prime})|\;\leqslant\;|B(\boldsymbol{x};r)|\;\leqslant\;\sum_{r^{\prime}=0}^{r}|A(\boldsymbol{x};r^{\prime})|$ (59)
and hence
$\max_{r^{\prime}\in\{0,\ldots,r\}}P_{q}(n;r^{\prime})\;\leqslant\;T_{q}(n;r)\;\leqslant\;\sum_{r^{\prime}=0}^{r}P_{q}(n;r^{\prime})\;\leqslant\;(r+1)\max_{r^{\prime}\in\{0,\ldots,r\}}P_{q}(n;r^{\prime}).$ (60)
It follows from (60) that $\beta_{q}(\rho)=\max_{\rho^{\prime}\in[0,\rho]}\alpha_{q}(\rho^{\prime})$, which is, by the fact that $\alpha_{q}(\rho)$ is concave and maximized for $\rho=\rho^{*}$, equal to the right-hand side of (58). $\blacksquare$
We now use the above result to obtain an upper bound on $\bar{\beta}_{q}(\rho)$. Note that, by Lemma 1, the domain of this function is $0,1]\cup\{\infty\}$.
Lemma 7
For any $q\geqslant 2$ and $\rho\in[0,1]$,
$\bar{\beta}_{q}(\rho)\;\leqslant\;\max_{\lambda\in[\frac{\rho}{3},\rho]}\left[2\lambda H\Big{(}\frac{\rho-\lambda}{2\lambda}\Big{)}+\beta_{q}(\lambda)\right],$ (61)
where $\beta_{q}(\cdot)$ is given in Lemma 6.
Proof
Lemma 4 implies that
$\bar{T}_{q}(n;r)$ $\;\leqslant\;\Big{(}\frac{2r}{3}+1\Big{)}\max_{v\in\{0,\ldots,2r/3\}}\binom{2(r-v)}{v}T_{q}(n;r-v)$
$\;=\;\Big{(}\frac{2r}{3}+1\Big{)}\max_{l\in\{r/3,\ldots,r\}}\binom{2l}{r-l}T_{q}(n;l).$ (62)
This, together with the fact that $\lim_{n\to\infty}\frac{1}{n}\log_{2}\!\binom{\gamma n}{\theta n}=\gamma H(\frac{\theta}{\gamma})$, gives (61). $\blacksquare$
4.3 Lower bound on $\bar{R}^{\star}_{q}(\delta)$
The following statement gives a lower bound on the achievable asymptotic rates of codes of relative distance $\delta$ (with respect to the distance function (4)).
Theorem 4.1
For any $q\geqslant 2$ and $\delta\in[0,1]$,
$\bar{R}^{\star}_{q}(\delta)\;\geqslant\;2\log_{2}(q)-\max_{\lambda\in[\frac{\delta}{3},\delta]}\left[2\lambda H\Big{(}\frac{\delta-\lambda}{2\lambda}\Big{)}+\beta_{q}(\lambda)\right],$ (63)
where the function $\beta_{q}(\cdot)$ is given in Lemma 6.
Proof
The result follows from the generalized Gilbert–Varshamov bound (44) and the upper bound on $\bar{\beta}_{q}(\cdot)$ given in Lemma 7. $\blacksquare$
Remark 3 (Lower bound for the binary case)
For $q=2$, a better lower bound on $\bar{R}^{\star}_{2}(\delta)$ is known [9, Section IV.C], as is the case in the regime of constant $t$ (see Remark 2). This is again due to the fact that the binary case seems to be rather special and admits a geometric interpretation in terms of packings in the Manhattan metric. $\blacklozenge$
5 Codes correcting all patterns of transpositions
It turns out that exponentially large codes of minimum distance $\infty$ exist in the model studied in this paper. In other words, the zero-error capacity of the transposition channel is positive, for any $q\geqslant 2$. For the special case $q=2$, it was shown in [21, Corollary 20] that $\bar{R}^{\star}_{2}(\infty)\geqslant\log_{2}(\lambda_{1})\approx 0.587$, where $\lambda_{1}$ is the largest root of the polynomial $p_{1}(x)=x^{7}-3x^{4}-2$. This was improved in [2, Theorem 3] to
$\bar{R}^{\star}_{2}(\infty)\;\geqslant\;\log_{2}(\lambda_{2})\;\approx\;0.643,$ (64)
where $\lambda_{2}\approx 1.561$ is the largest root of the polynomial $p_{2}(x)=x^{6}-2x^{3}-2x^{2}-2$. For $q$ divisible by 33, it was shown in [21, Theorem 21] that $\bar{R}^{\star}_{q}(\infty)\geqslant\log_{2}(\frac{q}{3})$. A construction that is valid for all $q$ and that improves upon [21, Theorem 21] was given in [2, Section IV], showing that
$\bar{R}^{\star}_{q}(\infty)\;\geqslant\;\frac{1}{2}\log_{2}\!\left(\left\lfloor\frac{q}{2}\right\rfloor\!\cdot\!\left\lceil\frac{q}{2}\right\rceil\right).$ (65)
In this section we provide a lower bound on $\bar{R}^{\star}_{q}(\infty)$ that is valid for arbitrary $q$, and that coincides with (64) for $q=2$ and improves upon (65) for $q\in\{3,4\}$. It is based on a code construction that represents a generalization of the construction from [2, Section III] for the binary case.
Note that any lower bound on $\bar{R}^{\star}_{q}(\infty)$ is automatically a lower bound on $\bar{R}^{\star}_{q}(\delta)$ for all $\delta\in[0,1]$, and is potentially better than (63) for $\delta>\delta_{0}$ (see Figure 1). In other words, for any $\delta\in[0,1]$, we can write
$\bar{R}^{\star}_{q}(\delta)\;\geqslant\;\max\!\left\{R^{\textsc{gv}}_{q}(\delta),\bar{R}^{\star}_{q}(\infty)\right\},$ (66)
where $R^{\textsc{gv}}_{q}(\delta)$ is the expression on the right-hand side of (63).
Construction:
Let $\{\mathcal{A}_{0},\mathcal{A}_{1}\}$ be a partition of the alphabet $\mathbb{Z}_{q}$, and
$\mathcal{P}\coloneqq\big{\{}(a,a,a),\,$ $(b,b,b),(a,b,b,b),(b,a,a,a),$
$(a,a,b,b,b,b),(b,b,a,a,a,a)\colon a\in\mathcal{A}_{0},b\in\mathcal{A}_{1}\big{\}}.$ (67)
Denote by $\mathcal{D}_{q}(n)$ the set of all strings of length $n$ that can be obtained as concatenations of the strings from $\mathcal{P}$.
Theorem 5.1
For every $q\geqslant 2$ and $n\geqslant 1$, we have $d_{\min}(\mathcal{D}_{q}(n))=\infty$, i.e., the code $\mathcal{D}_{q}(n)$ corrects all possible patterns of transpositions. Consequently,
$\bar{R}^{\star}_{q}(\infty)\;\geqslant\;\log_{2}(\lambda_{q}),$ (68)
where $\lambda_{q}$ is the unique positive root of the polynomial $p_{q}(x)=x^{6}-qx^{3}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil x^{2}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$.
Proof
For the first part of the statement, we need to show that an arbitrary codeword $\boldsymbol{x}\in\mathcal{D}_{q}(n)$ can be unambiguously recovered from the channel output $\boldsymbol{y}$ obtained after arbitrarily many transpositions have been applied to $\boldsymbol{x}$. Let $\boldsymbol{\tilde{x}}=(\tilde{x}_{1},\ldots,\tilde{x}_{n})$ denote the binary indicator string for $\boldsymbol{x}$ with respect to the partition $\{\mathcal{A}_{0},\mathcal{A}_{1}\}$, meaning that $\tilde{x}_{i}=j$ if and only if $x_{i}\in\mathcal{A}_{j}$, $j\in\{0,1\}$, and let $\widetilde{\mathcal{D}}(n)\coloneqq\{\boldsymbol{\tilde{x}}\colon\boldsymbol{x}\in\mathcal{D}_{q}(n)\}$. It was shown in [2, Theorem 3] that $\widetilde{\mathcal{D}}(n)$ is a binary code correcting all patterns of transpositions. We now have the following situation: the receiver can infer $\boldsymbol{\tilde{y}}$ from $\boldsymbol{y}$ (because it knows the sets $\mathcal{A}_{0},\mathcal{A}_{1}$), and then infer $\boldsymbol{\tilde{x}}$ from $\boldsymbol{\tilde{y}}$ (because $\boldsymbol{\tilde{x}}$ is a codeword of a transposition-correcting code), and what is left for it to do is infer $\boldsymbol{x}$ from $\boldsymbol{\tilde{x}}$. It is easy to see, by analyzing the possible concatenations of blocks from (67), that this can always be done. E.g., $(a,a,a,a^{\prime},a^{\prime},a^{\prime})$ may have resulted in $(a,a,a^{\prime},a,a^{\prime},a^{\prime})$ in the channel, $(a,a,a,a^{\prime},b,b,b)$ may have resulted in $(a,a,a^{\prime},a,b,b,b)$, etc. (here $a,a^{\prime}\in\mathcal{A}_{0}$, $b\in\mathcal{A}_{1}$), but such errors can easily be recognized and corrected because $(a,a,a^{\prime})$ is not a valid block. This means that the decoding process $\boldsymbol{y}\to\boldsymbol{\tilde{y}}\to\boldsymbol{\tilde{x}}\to\boldsymbol{x}$ is unambiguous, and hence the code $\mathcal{D}_{q}(n)$ corrects all patterns of transpositions.
It follows from the construction that the code size satisfies the recurrence relation
$|\mathcal{D}_{q}(n)|=$ $|\mathcal{A}_{0}|\cdot|\mathcal{D}_{q}(n-3)|+|\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-3)|$
$+2|\mathcal{A}_{0}||\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-4)|+2|\mathcal{A}_{0}||\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-6)|.$ (69)
This implies that $|\mathcal{D}_{q}(n)|=\lambda_{q}^{n+o(n)}$ as $n\to\infty$, where $\lambda_{q}$ is the unique positive root of the characteristic polynomial $p_{q}(x)=x^{6}-(|\mathcal{A}_{0}|+|\mathcal{A}_{1}|)x^{3}-2|\mathcal{A}_{0}||\mathcal{A}_{1}|x^{2}-2|\mathcal{A}_{0}||\mathcal{A}_{1}|$. Taking $|\mathcal{A}_{0}|=\lfloor\frac{q}{2}\rfloor$, $|\mathcal{A}_{1}|=\lceil\frac{q}{2}\rceil$, the second part of the statement follows. $\blacksquare$
We note that the code rate obtained in Theorem 5.1 is larger than that in (65) for $q\leqslant 4$, but not for $q>4$.
Using (66), (65), and Theorems 4.1 and 5.1, we have, for any $q\geqslant 2$ and any $\delta\in[0,1]$,
$\bar{R}^{\star}_{q}(\delta)\;\geqslant\;\max\!\left\{R^{\textsc{gv}}_{q}(\delta),\,\log_{2}(\lambda_{q}),\,\frac{1}{2}\log_{2}\!\left(\left\lfloor\frac{q}{2}\right\rfloor\!\cdot\!\left\lceil\frac{q}{2}\right\rceil\right)\right\},$ (70)
where $R^{\textsc{gv}}_{q}(\delta)$ is the expression on the right-hand side of (63), and $\lambda_{q}$ is the unique positive root of the polynomial $p_{q}(x)=x^{6}-qx^{3}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil x^{2}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$. The expression $\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$ simplifies to $\frac{q^{2}}{4}$ for even $q$, and $\frac{q^{2}-1}{4}$ for odd $q$.
The bound (70) is plotted in Figure 1 for the case $q=4$.
Figure 1: Lower bounds on the asymptotic rate of optimal transposition-correcting codes of relative distance $\delta$ over an alphabet of size $q=4$; see (70). The bound from Theorem 5.1 is better than that from Theorem 4.1 for $\delta>\delta_{0}\approx 0.343$.
6 Conclusion and further work
We have analyzed the problem of correcting transpositions of consecutive symbols in $q$-ary strings—types of errors that may occur in various contexts, e.g., as spelling errors, timing errors, inter-symbol interference errors (such as bit-shifts in magnetic storage devices), etc. As one of the main questions left open in this line of work we mention a construction of codes correcting $t>1$ transpositions that are asymptotically optimal in the sense that their cardinality scales as $\Theta(\frac{q^{n}}{n^{t}})$ when $n\to\infty$ (see (39)). In the linear regime $t=\tau n$, we have presented a lower bound on the achievable asymptotic code rates. Since the code space is nonuniform, obtaining good upper bounds is a nontrivial problem that is also left for future work, as are explicit constructions of asymptotically good (i.e., having positive rate) codes correcting a fraction of $\tau$ transpositions. For codes of infinite minimum distance, the problem of deriving the exact value of the maximum achievable rate (i.e., the zero-error capacity) is very interesting, and is still open for all alphabet sizes.
Finally, as mentioned in Section 1.3, there are several related models describing communication in the presence of synchronization and reordering errors that are potentially relevant for modern information transmission and storage systems and that are not yet fully understood, e.g., timing channels, non-binary channels with successive transpositions, permutation channels with bounded displacements, channels with a combination of transpositions and other types of errors, etc.
6.0.1 Acknowledgments.
M. Kovačević was supported by the Ministry of Science, Technological Development and Innovation of the Republic of Serbia (contract no. 451-03-137/2025-03/200156) and by the Faculty of Technical Sciences, University of Novi Sad, Serbia (project no. 01-50/295).