### v2

# Bounds on Codes Correcting Transpositions of Consecutive Symbols

Abstract
The problem of correcting transpositions (or swaps) of consecutive symbols in $q$-ary strings is studied. Lower bounds on asymptotically achievable rates of codes correcting $t=\tau n$ transpositions are derived. The first bound is obtained by analyzing the average cardinality of “transposition balls” and evaluating the appropriate version of the generalized Gilbert–Varshamov bound, while the second bound follows from a construction of codes correcting an arbitrary number of transpositions (i.e., zero-error codes). Asymptotic bounds on the cardinality of optimal codes correcting $t=\textrm{const}$ transpositions are also derived.
1 Introduction and preliminaries
Communication channels with various types of reordering and synchronization errors have received increased attention from the research community lately, for the most part due to their applications in modeling modern and somewhat unconventional information transmission and storage systems. In this paper we focus on a model where the transmitted strings are potentially altered by transpositions (or swaps) of consecutive symbols. These and similar kinds of impairments occur in various contexts, e.g., as spelling errors [7], timing and synchronization errors such as packet reordering in communication networks [3] or transpositions in molecular communications [14], intersymbol interference errors such as peak-shifts in magnetic recording channels [29] or charge drift in flash memories [2], genome rearrangement errors [6] which are of relevance for DNA-based data storage systems [9], etc.
In the remainder of this section we describe and discuss the model we are interested in, as well as another closely related model. In particular, we introduce a distance function that is appropriate for our scenario in the sense that it gives a characterization of the corresponding error correction problem (Section 1.1). We then analyze the effects of transpositions on $q$-ary strings and the cardinalities of appropriately defined balls (Section 2). We further derive lower bounds on the largest achievable asymptotic rate of codes correcting $t=\tau n$ transpositions. The first bound is a generalized Gilbert–Varshamov bound (Section 3.1). The second bound is a lower bound on the zero-error capacity of the transposition channel and follows from constructions of codes correcting all possible patterns of transpositions, i.e., zero-error codes. One such construction is presented herein and is shown to improve the best known constructions for $q=3,4$ (Section 3.2). Finally, we obtain asymptotic bounds on the cardinality of optimal codes correcting $t=\textrm{const}$ transpositions (Section 4).
1.1 Model description and basic definitions
Throughout the paper, $q$ is assumed to be an integer greater than or equal to 22, and the channel alphabet is denoted by $\mathbb{Z}_{q}=\{0,1,\ldots,q-1\}$. The possible inputs and outputs from the channel are the strings from $\mathbb{Z}_{q}^{n}$, where $n\geqslant 1$ is referred to as the block-length.
For a transmitted string $\boldsymbol{x}=(x_{1},\ldots,x_{n})$, we say that a transposition [^1] Transpositions of consecutive symbols are usually called adjacent transpositions in the literature. We refer to them simply as transpositions. No confusion should arise as this is the only kind of transpositions discussed in the paper. at location $k\in\{1,\ldots,n-1\}$ has occurred in the channel if the string obtained at the channel output is $\boldsymbol{y}=(y_{1},\ldots,y_{n})$, where $y_{k}=x_{k+1}$, $y_{k+1}=x_{k}$, and $y_{i}=x_{i}$ for all $i\notin\{k,k+1\}$, i.e., if the symbols $x_{k}$ and $x_{k+1}$ have been swapped. If $\boldsymbol{x}$ is altered by $t\geqslant 1$ transpositions, it is assumed that these transpositions are applied simultaneously to $\boldsymbol{x}$, and that therefore they affect disjoint pairs of symbols. In other words, if the transpositions occur at locations $k_{1},\ldots,k_{t}$, it is understood that $k_{j+1}\geqslant k_{j}+2$ for $j=1,\ldots,t-1$. Consequently, every symbol of $\boldsymbol{x}$ can be displaced by at most one position in the channel. Here is an example of a transmitted string $\boldsymbol{x}\in\mathbb{Z}_{4}^{10}$ and the received string $\boldsymbol{y}$ obtained after $t=3$ transpositions (the pairs of symbols that are being swapped are underlined):
[…]
$\hookrightarrow 10$ 1 $0302212\ =\ \boldsymbol{y}.$
We denote by $B(\boldsymbol{x};t)$ the set of all strings that can be obtained by applying at most $t$ transpositions, of the kind just described, to the string $\boldsymbol{x}$. It will be helpful to also introduce special notation $B^{\textnormal{e}}(\boldsymbol{x};t)$ for the set of all strings that can be obtained by applying exactly $t$ transpositions to $\boldsymbol{x}$ but cannot be obtained by applying fewer than $t$ transpositions to $\boldsymbol{x}$. The set $B(\boldsymbol{x};t)$ is therefore a disjoint union of the sets $B^{\textnormal{e}}(\boldsymbol{x};t^{\prime})$ over all $t^{\prime}\in\{0,1,\ldots,t\}$,
$B(\boldsymbol{x};t)=\bigsqcup_{t^{\prime}=0}^{t}B^{\textnormal{e}}(\boldsymbol{x};t^{\prime}).$ (2)
Note that $B^{\textnormal{e}}(\boldsymbol{x};t)=\emptyset$ for $n<2t$ because the largest possible number of disjoint transpositions in a string of length $n$ is $\lfloor n/2\rfloor$.
A nonempty subset of $\mathbb{Z}_{q}^{n}$ is called a code of length $n$; its elements are called codewords. A code $\mathcal{C}$ is said to be able to correct $t$ transpositions if the following condition holds:
$\forall\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}\qquad\boldsymbol{x}\neq\boldsymbol{x}^{\prime}\;\Rightarrow\;B(\boldsymbol{x};t)\cap B(\boldsymbol{x}^{\prime};t)=\emptyset.$ (3)
In words, any codeword $\boldsymbol{x}\in\mathcal{C}$, after being altered by up to $t$ transpositions in the channel, can be uniquely recovered from the received string (which belongs to $B(\boldsymbol{x};t)$) because no other codeword could have produced that string. Let $M^{\star}_{q}(n;t)$ denote the maximum cardinality of a code in $\mathbb{Z}_{q}^{n}$ correcting $t$ transpositions,
$M^{\star}_{q}(n;t)\coloneqq\max_{\mathcal{C}\subseteq\mathbb{Z}_{q}^{n},\,\mathcal{C}\,\textnormal{corrects $t$ transp.}}|\mathcal{C}|.$ (4)
We say $\mathcal{C}$ is able to correct an arbitrary number of transpositions (or that it is a zero-error code) if the relation (3) holds for every $t\geqslant 0$. Since the largest possible number of disjoint transpositions in a string of length $n$ is $\lfloor n/2\rfloor$, this is equivalent to writing
$\forall\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}\qquad\boldsymbol{x}\neq\boldsymbol{x}^{\prime}\;\Rightarrow\;B(\boldsymbol{x};\lfloor n/2\rfloor)\cap B(\boldsymbol{x}^{\prime};\lfloor n/2\rfloor)=\emptyset.$ (5)
In other words, the statements that a code corrects $\lfloor n/2\rfloor$ transpositions and that it corrects an arbitrary number of transpositions are equivalent.
We define the transposition distance between $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{q}^{n}$ by
$d(\boldsymbol{x},\boldsymbol{y})\coloneqq\min_{r,s\colon B(\boldsymbol{x};r)\cap B(\boldsymbol{y};s)\neq\emptyset}r+s,$ (6)
with the convention $\min\emptyset=\infty$. In words, we look at all the ways a common string $\boldsymbol{z}$ can be produced by applying transpositions to both $\boldsymbol{x}$ and $\boldsymbol{y}$, and among them we choose one requiring the smallest possible cumulative number of transpositions. That number is denoted by $d(\boldsymbol{x},\boldsymbol{y})$. If no common string can be produced by applying transpositions to $\boldsymbol{x}$ and $\boldsymbol{y}$, i.e., if $B(\boldsymbol{x};r)\cap B(\boldsymbol{y};s)=\emptyset$ for all $r,s$, we set $d(\boldsymbol{x},\boldsymbol{y})=\infty$. Again, since the largest possible number of disjoint transpositions in a string of length $n$ is $\lfloor n/2\rfloor$, we have
$d(\boldsymbol{x},\boldsymbol{y})=\infty\quad\Leftrightarrow\quad B(\boldsymbol{x};\lfloor n/2\rfloor)\cap B(\boldsymbol{y};\lfloor n/2\rfloor)=\emptyset.$ (7)
The ball of radius $t$ around $\boldsymbol{x}$ with respect to the distance $d(\cdot,\cdot)$ is denoted by
$\bar{B}(\boldsymbol{x};t)\coloneqq\{\boldsymbol{y}\colon d(\boldsymbol{x},\boldsymbol{y})\leqslant t\}.$ (8)
One can think of $\bar{B}(\boldsymbol{x};t)$ as the set of all strings that can be obtained through a two-step process: first apply $r$ transpositions to $\boldsymbol{x}$, and then apply $s$ transpositions to the resulting string, for any $r,s$ satisfying $r+s\leqslant t$. We note that $B(\boldsymbol{x};t)\subseteq\bar{B}(\boldsymbol{x};t)$, and that this containment relation is in most cases strict.
The function $d(\boldsymbol{x},\boldsymbol{y})$ is symmetric, non-negative, and equals zero if and only if $\boldsymbol{x}=\boldsymbol{y}$. However, it does not satisfy the triangle inequality and is therefore not a metric. E.g., for $\boldsymbol{x}=(1,0,0,0)$, $\boldsymbol{y}=(0,0,0,1)$, $\boldsymbol{z}=(0,0,1,0)$, we have $d(\boldsymbol{x},\boldsymbol{y})=\infty$, while $d(\boldsymbol{x},\boldsymbol{z})+d(\boldsymbol{z},\boldsymbol{y})=2+1=3$. Nonetheless, it is a convenient way of measuring distance between strings in the context of the problem currently being studied and gives a sufficient condition for the error-correction capability of codes.
Define the minimum distance of a code $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ in the usual way:
$d_{\min}(\mathcal{C})\coloneqq\min_{\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C},\;\boldsymbol{x}\neq\boldsymbol{x}^{\prime}}d(\boldsymbol{x},\boldsymbol{x}^{\prime}).$ (9)
Proposition 1
Let $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$, $\mathcal{C}\neq\emptyset$. (a) If $d_{\min}(\mathcal{C})>2t$, then $\mathcal{C}$ corrects $t$ transpositions; (b) $\mathcal{C}$ corrects an arbitrary number of transpositions if and only if $d_{\min}(\mathcal{C})=\infty$.
Proof
We first show the contrapositive of (a). If the code $\mathcal{C}$ is not able to correct $t$ transpositions, then, by (3), there exist distinct codewords $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}$ such that $B(\boldsymbol{x};t)\cap B(\boldsymbol{x}^{\prime};t)\neq\emptyset$. By (6) and (9), this implies that $d_{\min}(\mathcal{C})\leqslant d(\boldsymbol{x},\boldsymbol{x}^{\prime})\leqslant 2t$.
The statement (b) follows from (5) and (7). $\blacksquare$
The opposite statement of (a) is in general not true. For example, let $\mathcal{C}=\{\boldsymbol{x},\boldsymbol{y}\}\subseteq\mathbb{Z}_{2}^{10}$, where $\boldsymbol{x}=(1,0,1,0,0,0,1,0,1,0)$, $\boldsymbol{y}=(0,0,1,1,0,0,0,0,1,1)$. It is easy to see that the only common string that can be produced by applying transpositions to $\boldsymbol{x}$ and $\boldsymbol{y}$ is $\boldsymbol{z}=(0,1,0,1,0,0,0,1,0,1)$, and that the only way this can be done is by applying transpositions at locations $1,3,7,9$ to $\boldsymbol{x}$, and transpositions at locations $2,8$ to $\boldsymbol{y}$. This implies that $d(\boldsymbol{x},\boldsymbol{y})=4+2=6$, and also that $B(\boldsymbol{x};3)\cap B(\boldsymbol{y};3)=\emptyset$. In other words, the code $\mathcal{C}$ corrects $t=3$ transpositions, and yet $d_{\min}(\mathcal{C})\leqslant 2t$.
Remark 1(Transposition distance)
Another possible definition of “distance” between $\boldsymbol{x}$ and $\boldsymbol{y}$ that may seem natural in this context is the minimum number of transpositions needed to transform $\boldsymbol{x}$ into $\boldsymbol{y}$ (or, equivalently, $\boldsymbol{y}$ into $\boldsymbol{x}$), namely $d^{\prime}(\boldsymbol{x},\boldsymbol{y})\coloneqq\min\{t\colon\boldsymbol{y}\in B(\boldsymbol{x};t)\}$. However, this version of distance is not appropriate for the problem at hand because the corresponding analog of Proposition 1 does not hold. E.g., for $\boldsymbol{x}=(1,0,0)$, $\boldsymbol{y}=(0,0,1)$, we have $d^{\prime}(\boldsymbol{x},\boldsymbol{y})=\infty>2$, but $\{\boldsymbol{x},\boldsymbol{y}\}$ is not a code correcting one transposition because $\boldsymbol{z}=(0,1,0)$ can be obtained from both $\boldsymbol{x}$ and $\boldsymbol{y}$ through one transposition, i.e., $\boldsymbol{z}\in B(\boldsymbol{x};1)\cap B(\boldsymbol{y};1)$. $\blacklozenge$
1.2 A related model – successively performed transpositions
We briefly discuss here another model that is closely related to the one introduced in the previous subsection.
If the transmitted string is distorted by $t$ transpositions in the channel, one can imagine a scenario wherein these transpositions are performed in a successive manner, one after another, on pairs of symbols that are not necessarily disjoint. Here is an example with $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{4}^{10}$ and $t=3$:
[…]
In this model, pairs of symbols that are being swapped in two different steps may overlap, and thus one particular symbol of the transmitted string may end up up to $t$ positions away from its original position.
Let $B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$ denote the set of all strings that can be obtained by applying $t$ or fewer successive transpositions to the string $\boldsymbol{x}$.
The definition of distance appropriate for this model is the following:
$d^{\textnormal{{\tiny(s)}}}(\boldsymbol{x},\boldsymbol{y})\coloneqq\min\!\big\{t\colon\boldsymbol{y}\in B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)\big\},$ (11)
with the convention $\min\emptyset=\infty$. In words, $d^{\textnormal{{\tiny(s)}}}(\boldsymbol{x},\boldsymbol{y})$ is the smallest number of transpositions, performed in a successive manner, that can transform $\boldsymbol{x}$ into $\boldsymbol{y}$ (or, equivalently, $\boldsymbol{y}$ into $\boldsymbol{x}$). If $\boldsymbol{y}$ cannot be obtained from $\boldsymbol{x}$ by using any number of transpositions, we set $d(\boldsymbol{x},\boldsymbol{y})=\infty$. Unlike $d(\cdot,\cdot)$, $d^{\textnormal{{\tiny(s)}}}(\cdot,\cdot)$ is a metric [^2] This metric was analyzed on the space of permutations over $\{1,\ldots,n\}$ in, e.g., [20], but can be defined for arbitrary strings.. The set $B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$ defined above is a ball of radius $t$ around $\boldsymbol{x}$ with respect to this metric. Defining the minimum distance of a code $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ with respect to $d^{\textnormal{{\tiny(s)}}}$ in the usual way (see (9)), we have the following claim.
Proposition 2
A code $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ corrects $t$ transpositions (performed in a successive manner) if and only if $d^{\textnormal{{\tiny(s)}}}_{\min}(\mathcal{C})>2t$. $\blacksquare$
The model with successive transpositions is stronger than the one with simultaneous, disjoint transpositions, in the sense that all error patterns that can be realized in the latter can also be realized in the former, i.e., $B(\boldsymbol{x};t)\subseteq B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$. (In fact, the stronger statement $\bar{B}(\boldsymbol{x};t)\subseteq B^{\textnormal{{\tiny(s)}}}(\boldsymbol{x};t)$ is also true.) Consequently: 1.) any code construction valid for the successive transposition model is also valid for the disjoint transposition model, 2.) any lower bound on the cardinality of optimal codes for the successive transposition model is also valid for the disjoint transposition model, and 3.) any upper bound on the cardinality of optimal codes in the disjoint transposition model is also valid in the successive transposition model. Therefore, the upper bound on the cardinality of optimal codes correcting $t$ disjoint transpositions presented in Theorem 4.1 is automatically valid for the stronger model with successive transpositions as well.
1.3 Transpositions vs. substitutions
For $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$, let $\boldsymbol{\hat{x}}\in\mathbb{Z}_{q}^{n}$ be the string defined by $\hat{x}_{i}=\sum_{j=1}^{i}x_{j}$, $i=1,\ldots,n$, where addition is performed modulo $q$. Equivalently, $x_{i}=\hat{x}_{i}-\hat{x}_{i-1}$, $i=1,\ldots,n$, where $x_{0}=0$ and subtraction is performed modulo $q$. Then it is easy to see that any $t$ transpositions in $\boldsymbol{x}$ become $t$ substitutions in $\boldsymbol{\hat{x}}$. This is true not only for disjoint but for successive transpositions as well. It follows that, if $\widehat{\mathcal{C}}\subseteq\mathbb{Z}_{q}^{n}$ is a code correcting $t$ substitutions, i.e., a code having minimum Hamming distance larger than $2t$, then $\mathcal{C}=\{\boldsymbol{x}\colon\boldsymbol{\hat{x}}\in\widehat{\mathcal{C}}\}$ is a code correcting $t$ transpositions. Consequently, any lower bound on the cardinality of optimal codes correcting $t$ substitutions is automatically a lower bound on the cardinality of optimal codes correcting $t$ transpositions.
Since the transposition model is in this sense weaker than the substitution model, we expect to be able to achieve higher code rates in the former. This is indeed demonstrated in Section 3.1 and illustrated in Figure 1.
1.4 Prior work
Asymptotically achievable rates of codes correcting $t=\tau n$ transpositions have been studied previously only in a few special cases. In the binary case ($q=2$), transpositions are also known as bit-shifts or peak-shifts in the literature and are of relevance for modeling errors in magnetic recording devices. Lower bounds on achievable rates of binary codes correcting a fraction of $\tau$ (successively performed) bit-shifts were derived in [19, 12]. Over general alphabets, codes correcting all possible patterns of transpositions, for the model studied herein as well as for the more general model where each symbol can be displaced by up to $\ell$ positions, were studied in [26, 4]. These code constructions imply lower bounds on the zero-error capacity of the transposition channel or, equivalently, on the largest achievable rate of codes correcting a fraction of $\tau=\frac{1}{2}$ transpositions. In the present paper we provide lower bounds on achievable rates valid for any $q\geqslant 2$ and any $\tau\in[0,\frac{1}{2}]$.
We mention in this context also the works studying zero-error codes for models that are not equivalent to the model analyzed here, but are closely related to it, namely the binary bit-shift channel (with an additional assumption that each 11 can be shifted by at most $\ell$ positions) [29, 24] and several types of timing channels [22, 23].
As for codes correcting a fixed number of transpositions, constructions and bounds for the binary case were presented in [21] (for the stronger model with successively performed transpositions/bit-shifts). Constructions of binary codes correcting a single transposition appear in some earlier works as well, e.g., in [25, 27]. In the present paper we obtain asymptotic bounds on the cardinality of optimal $q$-ary codes correcting $t$ transpositions, for any $q\geqslant 2$ and $t\geqslant 1$.
Codes for models that include other types of errors in addition to transpositions, e.g., deletions and/or substitutions, were studied in [1, 5, 7, 9, 10, 16, 17, 28, 32, 33].
Finally, let us mention that transposition errors are also of interest in another scenario where the stored data is represented by a permutation on $\{1,\ldots,n\}$. Namely, transposition-correcting codes in the space of all permutations occur naturally in rank modulation schemes for flash memories; see, e.g., [2].
1.5 Notation
By a run in a string $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$ we mean a substring of identical symbols that is delimited on both sides either by a different symbol or by the end of the string. The number of runs in $\boldsymbol{x}$ is denoted by $\operatorname{run}(\boldsymbol{x})$. For example, the string $\boldsymbol{x}=(0,3,3,3,0,0,1,2,2)\in\mathbb{Z}_{4}^{9}$ has $\operatorname{run}(\boldsymbol{x})=5$ runs.
$H_{q}(x)=-x\log_{q}x-(1-x)\log_{q}(1-x)+x\log_{q}(q-1)$ is the $q$-ary entropy function. As usual, we write simply $H(x)$ for $H_{2}(x)$.
In the asymptotic analysis in Section 4 we will need the following notation. For two nonnegative sequences $(a_{n})$, $(b_{n})$:
  • $\bullet$
    $a_{n}\sim b_{n}$ means $\lim_{n\to\infty}\frac{a_{n}}{b_{n}}=1$,
  • $\bullet$
    $a_{n}\lesssim b_{n}$ means $\limsup_{n\to\infty}\frac{a_{n}}{b_{n}}\leqslant 1$,
  • $\bullet$
    $a_{n}=o(b_{n})$ means $\lim_{n\to\infty}\frac{a_{n}}{b_{n}}=0$,
  • $\bullet$
    $a_{n}={\mathcal{O}}(b_{n})$ means $\limsup_{n\to\infty}\frac{a_{n}}{b_{n}}<\infty$,
  • $\bullet$
    $a_{n}=\Omega(b_{n})$ means $\liminf_{n\to\infty}\frac{a_{n}}{b_{n}}>0$, or equivalently $b_{n}={\mathcal{O}}(a_{n})$,
  • $\bullet$
    $a_{n}=\Theta(b_{n})$ means $0<\liminf_{n\to\infty}\frac{a_{n}}{b_{n}}\leqslant\limsup_{n\to\infty}\frac{a_{n}}{b_{n}}<\infty$.
2 Properties of transpositions and transposition distance
In this section we demonstrate several facts about the transposition distance and the sets $B(\boldsymbol{x};r)$ and $\bar{B}(\boldsymbol{x};r)$ that, in addition to being of interest in their own right, will be needed for the derivation of the bounds in subsequent sections.
Proposition 3
For every $q\geqslant 2$ and $n\geqslant 1$, the maximum distance between two strings $\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{q}^{n}$ such that $d(\boldsymbol{x},\boldsymbol{y})<\infty$ is $\max_{\boldsymbol{x},\boldsymbol{y}\in\mathbb{Z}_{q}^{n},\,d(\boldsymbol{x},\boldsymbol{y})<\infty}d(\boldsymbol{x},\boldsymbol{y})=n-1$.
Proof
Suppose a string $\boldsymbol{z}$ can be obtained by applying transpositions at locations $k_{1},\ldots,k_{t}$ to the string $\boldsymbol{x}$, and also by applying transpositions at locations $\ell_{1},\ldots,\ell_{s}$ to the string $\boldsymbol{y}$. If $k_{i}=\ell_{j}$, for some $i\in\{1,\ldots,t\}$, $j\in\{1,\ldots,s\}$, then $(z_{k_{i}},z_{k_{i}+1})=(x_{k_{i}+1},x_{k_{i}})=(y_{k_{i}+1},y_{k_{i}})$, and hence a common string $\boldsymbol{z}^{\prime}$ (obtained by applying a transposition at location $k_{i}$ to $\boldsymbol{z}$) can be obtained by applying t1t-1 transpositions (at locations $k_{1},\ldots,k_{i-1},k_{i+1},\ldots k_{t}$) to $\boldsymbol{x}$ and s1s-1 transpositions (at locations $\ell_{1},\ldots,\ell_{j-1},\ell_{j+1},\ldots\ell_{s}$) to $\boldsymbol{y}$. Therefore, when computing $d(\boldsymbol{x},\boldsymbol{y})$, we may, without loss of generality, assume that the sets of transposition locations $\{k_{1},\ldots,k_{t}\}$ and $\{\ell_{1},\ldots,\ell_{s}\}$ are disjoint. Since $k_{t},\ell_{s}\leqslant n-1$, this implies that $t+s\leqslant n-1$, and hence that $d(\boldsymbol{x},\boldsymbol{y})\leqslant n-1$. That this upper bound is attained is shown by the following examples. For even $n$, take $\boldsymbol{x}$ to be the alternating string $(1,0,1,0,\ldots)$ of length $n$, and $\boldsymbol{y}$ the concatenation of the string $(0)$, the alternating string $(0,1,0,1,\ldots)$ of length n2n-2, and the string $(1)$. E.g., for $n=6$, $\boldsymbol{x}=(1,0,1,0,1,0)$ and $\boldsymbol{y}=(0,0,1,0,1,1)$. Then the only common string that $\boldsymbol{x}$ and $\boldsymbol{y}$ can produce is the alternating string $\boldsymbol{z}=(0,1,0,1,\ldots)$, obtained by applying $n/2$ transpositions at locations $1,3,\ldots,n-1$ to $\boldsymbol{x}$, and n/21n/2-1 transpositions at locations $2,4,\ldots,n-2$ to $\boldsymbol{y}$. Similarly, for odd $n$, take $\boldsymbol{x}$ to be the concatenation of the alternating string $(1,0,1,0,\ldots)$ of length n1n-1 and the string $(0)$, and $\boldsymbol{y}$ the concatenation of the string $(0)$ and the alternating string $(0,1,0,1,\ldots)$. E.g., for $n=5$, $\boldsymbol{x}=(1,0,1,0,0)$ and $\boldsymbol{y}=(0,0,1,0,1)$. $\blacksquare$
Of particular interest for the present paper are the cardinalities of the sets $B(\boldsymbol{x};r)$ and $\bar{B}(\boldsymbol{x};r)$. Obtaining the exact characterization of these quantities appears to be difficult, but one can derive good bounds.
It is easy to see that
$|B(\boldsymbol{x};1)|=|B^{\textnormal{e}}(\boldsymbol{x};1)|+1=\operatorname{run}(\boldsymbol{x}).$ (12)
This is because only the transpositions occurring at the boundaries of consecutive runs in $\boldsymbol{x}$ will actually alter $\boldsymbol{x}$, and hence there are $\operatorname{run}(\boldsymbol{x})-1$ strings different from $\boldsymbol{x}$ that can be obtained in this way. Accounting for $\boldsymbol{x}$ itself, (12) follows. For general $r$, we have the bounds on $|B^{\textnormal{e}}(\boldsymbol{x};r)|$ stated in Proposition 4. Bounds on $|B(\boldsymbol{x};r)|$ can then be obtained from
$|B(\boldsymbol{x};r)|\;=\;\sum_{r^{\prime}=0}^{r}|B^{\textnormal{e}}(\boldsymbol{x};r^{\prime})|,$ (13)
which follows from (2).
Proposition 4
For every $q\geqslant 2$, $r\geqslant 1$, $n\geqslant 2r$, and $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$,
$\sum_{u=0}^{r}\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor}{u}\binom{\lceil\operatorname{run}(\boldsymbol{x})/2\rceil-2u-1}{r-u}\;\leqslant\;|B^{\textnormal{e}}(\boldsymbol{x};r)|\;\leqslant\;\binom{n-r}{r}$. (14)
Proof
To derive a lower bound on the cardinality of $B^{\textnormal{e}}(\boldsymbol{x};r)$, consider the odd-numbered runs (first, third, etc.) in $\boldsymbol{x}$. Every transposition in $\boldsymbol{x}$ occurs either at the left or at the right boundary of an odd-numbered run, and so every pattern of exactly $r$ transpositions can be described by assuming that $u$ transpositions occur at the right boundary of such a run in the first phase (i.e., we select $u$ odd-numbered runs and swap their right-most symbols with the left-most symbols of the even-numbered runs immediately to their right), after which rur-u transpositions occur at the left boundary of such a run in the second phase, for some $u=0,1,\ldots,r$. The former can happen in $\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor}{u}$ ways, and the latter in at least $\binom{\lceil\operatorname{run}(\boldsymbol{x})/2\rceil-2u-1}{r-u}$ ways, implying the left-hand inequality in (14). (In the second phase we exclude the first run, because it has no boundary on the left. Also, if the $i$’th run was selected for the transposition in the first phase, we exclude in the second phase both the $i$’th run, because it is possible that it is of length 11 in which case its left and right boundaries coincide, and the $(i+2)$’th run, because it is possible that the $(i+1)$’th run in between them was of length 11. Hence the 2u1-2u-1 term.)
To prove the stated upper bound on the cardinality of $B^{\textnormal{e}}(\boldsymbol{x};r)$, we need to show that there are at most $\binom{n-r}{r}$ strings that can be obtained by applying exactly $r$ transpositions to $\boldsymbol{x}$. Since the transposition locations $k_{1},\ldots,k_{r}$ belong to $\{1,\ldots,n-1\}$ and satisfy $k_{j+1}>k_{j}+1$, every set of $r$ transposition locations corresponds uniquely to a binary string with $r$ 11’s and nrn-r 0’s such that there is at least one 0 after every 11. The number of such strings can be found by counting the number of ways to distribute n2rn-2r 0’s in $r+1$ blocks (where the blocks are separated by pairs of symbols $1,0$), which is $\binom{(n-2r)+(r+1)-1}{(r+1)-1}=\binom{n-r}{r}$. $\blacksquare$
Proposition 5
For every $q\geqslant 2$, $n\geqslant 1$, $r\geqslant 1$, and $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$,
$|B(\boldsymbol{x};r)|\;\leqslant\;|\bar{B}(\boldsymbol{x};r)|\;\leqslant\;\sum_{u=\lceil r/3\rceil}^{\min\{r,\lfloor n/2\rfloor\}}\binom{2u}{r-u}|B(\boldsymbol{x};u)|.$ (15)
Proof
The left-hand inequality is clear from $B(\boldsymbol{x};r)\subseteq\bar{B}(\boldsymbol{x};r)$.
To show the right-hand inequality, think of $\bar{B}(\boldsymbol{x};r)$ as the set of all strings that can be obtained by first applying up to $u$ transpositions to $\boldsymbol{x}$, resulting in a string $\boldsymbol{y}$, and then applying $v$ transpositions to $\boldsymbol{y}$, where $u+v=r$.
Taking $u=r$ in the first step (and hence $v=0$ in the second), we get exactly $B(\boldsymbol{x};r)$, i.e., the set of strings that can be obtained from $\boldsymbol{x}$ with at most $r$ transpositions. This is the summand $u=r$ in (15).
Consider now the case u=r1u=r-1, and let $\boldsymbol{y}\in B(\boldsymbol{x};r-1)$ be a string obtained from $\boldsymbol{x}$ in the first step. In the second step we are allowed to choose $v=1$ pair of consecutive symbols in $\boldsymbol{y}$ to swap. If this pair is disjoint from all the pairs chosen in the first step, then the resulting string $\boldsymbol{z}$ can equivalently be obtained directly from $\boldsymbol{x}$ via up to $r$ disjoint transpositions. In other words, $\boldsymbol{z}\in B(\boldsymbol{x};r)$ and this string was therefore already accounted for in the first case described in the previous paragraph. Also, if the pair chosen in the second step is identical to a pair chosen in the first step, then the transposition from the second step will annul the transposition from the first step and, again, $\boldsymbol{z}\in B(\boldsymbol{x};r)$, i.e., we are back to the previous case. Therefore, in the second step we can, without loss of generality, consider only pairs of symbols that partially overlap with a pair chosen in the first step, either on its left or on its right symbol. (E.g., suppose that $(a,b,c,d)$ is a substring of $\boldsymbol{x}$, and that the pair $(b,c)$ was chosen in the first step, so that the corresponding substring of $\boldsymbol{y}$ is $(a,c,b,d)$. Then the pairs in $\boldsymbol{y}$ that partially overlap with the pair chosen in the first step, one of which we can choose in the second step, are $(a,c)$ and $(b,d)$.) Since at most $u$ pairs were chosen in the first step, there are at most $2u$ pairs to select one from in the second step. In conclusion, there are $|B(\boldsymbol{x};u)|$ different strings $\boldsymbol{y}$ that can be obtained in the first step, and for each such $\boldsymbol{y}$ there are at most $\binom{2u}{1}=2u$ different strings $\boldsymbol{z}$ that can be obtained in the second step.
The analysis in the general case is the same: there are $|B(\boldsymbol{x};u)|$ different strings $\boldsymbol{y}$ that can be obtained in the first step, and for each such $\boldsymbol{y}$ there are at most $\binom{2u}{v}$ different strings $\boldsymbol{z}$ that can be obtained in the second step.
Finally, note that, by the nature of transposition errors, $u\leqslant\lfloor n/2\rfloor$. Hence the upper bound on the summation index in (15). $\blacksquare$
We next define and analyze the average cardinalities of the sets $B^{\textnormal{e}}(\boldsymbol{x};r)$, $B(\boldsymbol{x};r)$, and $\bar{B}(\boldsymbol{x};r)$ over the whole space $\mathbb{Z}_{q}^{n}$. Let
$T^{\textnormal{e}}_{q}(n;r)$ $\coloneqq\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|B^{\textnormal{e}}(\boldsymbol{x};r)|,$ (16)
$T_{q}(n;r)$ $\coloneqq\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|B(\boldsymbol{x};r)|,$
$\bar{T}_{q}(n;r)$ $\coloneqq\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};r)|,$
so that the average value of $|B^{\textnormal{e}}(\boldsymbol{x};r)|$ is $\frac{1}{q^{n}}T^{\textnormal{e}}_{q}(n;r)$, and similarly for the others. It follows from (13), (15), and (16) that
$T_{q}(n;r)\;=\;\sum_{r^{\prime}=0}^{r}T^{\textnormal{e}}_{q}(n;r^{\prime})$ (17)
and
${T}_{q}(n;r)\;\leqslant\;\bar{T}_{q}(n;r)\;\leqslant\;\sum_{u=\lceil r/3\rceil}^{\min\{r,\lfloor n/2\rfloor\}}\binom{2u}{r-u}T_{q}(n;u).$ (18)
Proposition 6
For every $q\geqslant 2$, $r\geqslant 0$, and $n\geqslant 2r$,
$T^{\textnormal{e}}_{q}(n;r)=\binom{n-r}{r}q^{n-r}(q-1)^{r}.$ (19)
Proof
We first derive the generating function of the bivariate sequence $(T^{\textnormal{e}}_{q}(n;r))_{n,r}$.
Consider the two cases:
  • $\bullet$
    If the last two symbols of $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$ are different, i.e., $\boldsymbol{x}=\boldsymbol{x}^{\prime}yy^{\prime}$ with $\boldsymbol{x}^{\prime}\in\mathbb{Z}_{q}^{n-2}$, $^{\prime}\in\mathbb{Z}_{q}, yyy\neq y^{\prime}$, then
    $|B^{\textnormal{e}}(\boldsymbol{x};r)|=|B^{\textnormal{e}}(\boldsymbol{x}^{\prime}y;r)|+|B^{\textnormal{e}}(\boldsymbol{x}^{\prime};r-1)|.$ (20a)
    The first term corresponds to the case when all the transpositions occur in the substring $\boldsymbol{x}^{\prime}y$, and the second term to the case when r1r-1 transpositions occur in the substring $\boldsymbol{x}^{\prime}$ and one transposition occurs in $yy^{\prime}$.
  • $\bullet$
    If the last two symbols of $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$ are the same, i.e., $\boldsymbol{x}=\boldsymbol{x}^{\prime}yy$, then
    $|B^{\textnormal{e}}(\boldsymbol{x};r)|=|B^{\textnormal{e}}(\boldsymbol{x}^{\prime}y;r)|.$ (20b)
We then have
$T^{\textnormal{e}}_{q}(n;r)$ $=\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|B^{\textnormal{e}}(\boldsymbol{x};r)|$
$=\sum_{\boldsymbol{x}^{\prime}y\in\mathbb{Z}_{q}^{n-1}}\sum_{y^{\prime}\in\mathbb{Z}_{q}\setminus\{y\}}|B^{\textnormal{e}}(\boldsymbol{x}^{\prime}yy^{\prime};r)|+\sum_{\boldsymbol{x}^{\prime}yy\in\mathbb{Z}_{q}^{n}}|B^{\textnormal{e}}(\boldsymbol{x}^{\prime}yy;r)|$
$=\sum_{\boldsymbol{x}^{\prime}y\in\mathbb{Z}_{q}^{n-1}}\sum_{y^{\prime}\in\mathbb{Z}_{q}\setminus\{y\}}(|B^{\textnormal{e}}(\boldsymbol{x}^{\prime}y;r)|+|B^{\textnormal{e}}(\boldsymbol{x}^{\prime};r-1)|)+\sum_{\boldsymbol{x}^{\prime}y\in\mathbb{Z}_{q}^{n-1}}|B^{\textnormal{e}}(\boldsymbol{x}^{\prime}y;r)|$
$=(q-1)T^{\textnormal{e}}_{q}(n-1;r)+q(q-1)T^{\textnormal{e}}_{q}(n-2;r-1)+T^{\textnormal{e}}_{q}(n-1;r)$
$=qT^{\textnormal{e}}_{q}(n-1;r)+q(q-1)T^{\textnormal{e}}_{q}(n-2;r-1).$ (21)
From the definition of $B^{\textnormal{e}}(\boldsymbol{x};r)$ we can also find the initial conditions of the recurrence relation (21):
$T^{\textnormal{e}}_{q}(0;0)=1,\quad T^{\textnormal{e}}_{q}(0;r>0)=0,\quad T^{\textnormal{e}}_{q}(1;0)=q,\quad T^{\textnormal{e}}_{q}(1;r>0)=0.$ (22)
It is now straightforward to derive the generating function of $(T^{\textnormal{e}}_{q}(n;r))_{n,r}$ by using the recurrence relation (21) and the initial conditions (22),
$F^{\textnormal{e}}_{q}(z,w)\coloneqq\sum_{n,r\geqslant 0}T^{\textnormal{e}}_{q}(n;r)z^{n}w^{r}$ $=\frac{1}{1-qz-q(q-1)z^{2}w}$ (23)
$=\frac{1}{1-qz}\sum_{r=0}^{\infty}\left(\frac{q(q-1)z^{2}}{1-qz}\right)^{r}w^{r}.$
From (23) we obtain
$F^{\textnormal{e}}_{q}(z,w)$ $=\frac{(q(q-1)z^{2})^{r}}{(1-qz)^{r+1}}$ (24)
$=q^{r}(q-1)^{r}z^{2r}\sum_{m=0}^{\infty}\binom{r+m}{r}q^{m}z^{m}$
$=(q-1)^{r}\sum_{k=2r}^{\infty}\binom{k-r}{r}q^{k-r}z^{k},$
and hence, for $n\geqslant 2r$,
$T^{\textnormal{e}}_{q}(n;r)=[z^{n}w^{r}]F^{\textnormal{e}}_{q}(z,w)=\binom{n-r}{r}q^{n-r}(q-1)^{r},$ (25)
which was to be shown. $\blacksquare$
3 Achievable rates of codes correcting $t\!=\!\tau n$ transpositions
Let $R_{q}^{\star}(\tau)$ be the largest asymptotic rate of codes in $\mathbb{Z}_{q}^{n}$ correcting a fraction of $\tau\in[0,\frac{1}{2}]$ transpositions, expressed in bits per symbol,
${R}_{q}^{\star}(\tau)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}{M}_{q}^{\star}(n;\lfloor\tau n\rfloor).$ (26)
In this section we derive two lower bounds on the function $R_{q}^{\star}(\tau)$ (Theorems 3.1 and 3.2 ahead).
3.1 Generalized Gilbert–Varshamov bound
The standard Gilbert–Varshamov bound states that the cardinality of an optimal code of minimum distance $d$ is lower-bounded by the ratio of the cardinality of the code space and the cardinality of a ball of radius d1d-1 [11, 31]. The bound was later generalized [18, 13] to include nonuniform spaces in which the cardinality of a ball of given radius depends on its center. Namely, it was proved in [13] that, in such cases, the lower bound still holds if one puts in the denominator the average cardinality of a ball of radius d1d-1.
For $q\geqslant 2$ and $\rho\in[0,\frac{1}{2}]$, define
$\beta^{\textnormal{e}}_{q}(\rho)=(1-\rho)H\Big(\frac{\rho}{1-\rho}\Big)+(1-\rho)\log_{2}(q)+\rho\log_{2}(q-1),$ (27)
where $H(\cdot)$ is the binary entropy function. It is easy to check, by computing its derivatives, that the function $\beta^{\textnormal{e}}_{q}(\rho)$ is concave and maximized for
$\rho^{*}=\frac{1}{2}\left(1-\sqrt{\frac{q}{5q-4}}\right).$ (28)
Further, for $q\geqslant 2$ and $\rho\in[0,\frac{1}{2}]$, define
$\beta_{q}(\rho)=\max_{0\leqslant\rho^{\prime}\leqslant\rho}\beta^{\textnormal{e}}_{q}(\rho^{\prime})=\begin{cases}\beta^{\textnormal{e}}_{q}(\rho),&\rho\leqslant\rho^{*}\\ \beta^{\textnormal{e}}_{q}(\rho^{*}),&\rho\geqslant\rho^{*}\end{cases}.$ (29)
Theorem 3.1
For any $q\geqslant 2$ and $\tau\in[0,\frac{1}{2}]$,
${R}^{\star}_{q}(\tau)\;\geqslant\;2\log_{2}(q)-\max_{\frac{2\tau}{3}\leqslant\lambda\leqslant\min\{2\tau,\frac{1}{2}\}]}\left[2\lambda H\Big(\frac{2\tau-\lambda}{2\lambda}\Big)+\beta_{q}(\lambda)\right],$ (30)
where the function $\beta_{q}(\cdot)$ is defined in (29).
Proof
As noted in Section 1.1, the distance function $d(\cdot,\cdot)$ is not a metric, and hence $\bar{B}(\boldsymbol{x};r)$ is not a metric ball. Nonetheless, the same argument used to obtain the Gilbert–Varshamov bound in classical settings [11, 31, 13] can be applied in our setting as well. Namely, for any given $t$, by using a greedy procedure one is guaranteed to find a collection of strings $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ satisfying the following two conditions:
  • 1.)
    $d(\boldsymbol{x},\boldsymbol{x}^{\prime})>2t$ (i.e., $\boldsymbol{x}^{\prime}\notin\bar{B}(\boldsymbol{x};2t)$, see (8)) for any two distinct $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}$,
  • 2.)
    $\bigcup_{\boldsymbol{x}\in\mathcal{C}}\bar{B}(\boldsymbol{x};2t)\supseteq\mathbb{Z}_{q}^{n}$.
The first condition implies, by Proposition 1, that $\mathcal{C}$ is a code correcting $t$ transpositions. The second condition, stating that balls of radius $2t$ around the codewords cover the whole space, implies that $\sum_{\boldsymbol{x}\in\mathcal{C}}|\bar{B}(\boldsymbol{x};2t)|\geqslant q^{n}$. Therefore, it follows from the argument in [13] that
${M}_{q}^{\star}(n;t)\;\geqslant\;\frac{q^{n}}{\frac{1}{q^{n}}\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};2t)|}\;=\;\frac{q^{2n}}{\bar{T}_{q}(n;2t)},$ (31)
where $\bar{T}_{q}(n;r)$ is the “total ball size” of radius $r$ (see (16)).
If we denote the exponential growth rate (to base 22) of $\bar{T}_{q}(n;\lfloor\rho n\rfloor)$ by
$\bar{\beta}_{q}(\rho)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}\bar{T}_{q}(n;\lfloor\rho n\rfloor),$ (32)
then (31) implies that
${R}^{\star}_{q}(\tau)\;\geqslant\;2\log_{2}(q)-\bar{\beta}_{q}(2\tau).$ (33)
Therefore, an upper bound on the function $\bar{\beta}_{q}(\cdot)$ would imply a lower bound on the achievable asymptotic rates of transposition-correcting codes.
From Proposition 6 and the fact that $\lim_{n\to\infty}\frac{1}{n}\log_{2}\!\binom{\alpha n}{\gamma n}=\alpha H(\frac{\gamma}{\alpha})$ one can obtain the exponential growth rate of $T^{\textnormal{e}}_{q}(n;\lfloor\rho n\rfloor)$, namely
$\lim_{n\to\infty}\frac{1}{n}\log_{2}T^{\textnormal{e}}_{q}(n;\lfloor\rho n\rfloor)=\beta^{\textnormal{e}}_{q}(\rho),$ (34)
where $\beta^{\textnormal{e}}_{q}(\rho)$ is given by (27). Note that, since the maximum number of disjoint transpositions is $\lfloor n/2\rfloor$, the domain of this function is $0,\frac{1}{2}]$. Then from (17) one obtains the exponential growth rate of $T_{q}(n;\lfloor\rho n\rfloor)$, namely
$\lim_{n\to\infty}\frac{1}{n}\log_{2}T_{q}(n;\lfloor\rho n\rfloor)=\beta_{q}(\rho),$ (35)
where $\beta_{q}(\rho)$ is given by (29). Finally, Proposition 5 implies that
$\bar{T}_{q}(n;r)\;\leqslant\;\Big(\frac{2r}{3}+1\Big)\max_{\lceil\frac{r}{3}\rceil\leqslant u\leqslant\min\{r,\lfloor\frac{n}{2}\rfloor\}\}}\binom{2u}{r-u}T_{q}(n;u),$ (36)
which further implies that, for any $q\geqslant 2$ and $\rho\in[0,1]$,
$\bar{\beta}_{q}(\rho)\;\leqslant\;\max_{\frac{\rho}{3}\leqslant\lambda\leqslant\min\{\rho,\frac{1}{2}\}}\left[2\lambda H\Big(\frac{\rho-\lambda}{2\lambda}\Big)+\beta_{q}(\lambda)\right].$ (37)
The theorem follows from (33) and (37). $\blacksquare$
The bound from Theorem 3.1 is plotted in Figure 1 for an alphabet of size $q=4$. For comparison, we also plot the bound
${R}^{\star}_{q}(\tau)\;\geqslant\;\begin{cases}\log_{2}(q)-2\tau\log_{2}(q-1)-H(2\tau),&0\leqslant\tau\leqslant\frac{q-1}{2q}\\ 0,&\tau>\frac{q-1}{2q}\end{cases}.$ (38)
The expression on the right-hand side of (38) is the familiar Gilbert–Varshamov bound for codes correcting $\tau n$ substitutions, i.e., codes of minimum Hamming distance $2\tau n$ [15, Section 2.10.6]. Recall from Section 1.3 that any code correcting $t$ substitutions can be used to construct a code of the same cardinality correcting $t$ transpositions.
Remark 2(Lower bound for the binary case)
For $q=2$, a lower bound on ${R}^{\star}_{2}(\tau)$ better than (30) is known [12, Section IV.C]. This is due to the fact that binary codes correcting transposition (a.k.a. bit-shift) errors can be interpreted as packings in the Manhattan metric [21], which in particular leads to a finer estimate of $\bar{T}_{2}(n;r)$. Unfortunately, the mentioned geometric interpretation fails when $q>2$ and thus the corresponding analysis and the constructions via dense packings cannot be generalized to larger alphabets in any obvious way. $\blacklozenge$
3.2 Codes correcting an arbitrary number of transpositions and the zero-error bound
It turns out that exponentially large codes of minimum distance $\infty$ exist in the model studied in this paper. In other words, the zero-error capacity of the transposition channel – the largest asymptotic rate of codes correcting an arbitrary number of transpositions – is positive, for any $q\geqslant 2$. Note that, by (7), Proposition 1(a), and (26), this rate is precisely ${R}^{\star}_{q}(1/2)$.
For the special case $q=2$, it was shown in [26, Corollary 20] that ${R}^{\star}_{2}(1/2)\geqslant\log_{2}(\xi_{1})\approx 0.587$, where $\xi_{1}$ is the largest root of the polynomial $p_{1}(x)=x^{7}-3x^{4}-2$. This was improved in [4, Theorem 3] to
${R}^{\star}_{2}(1/2)\;\geqslant\;\log_{2}(\xi_{2})\;\approx\;0.643,$ (39)
where $\xi_{2}\approx 1.561$ is the largest root of the polynomial $p_{2}(x)=x^{6}-2x^{3}-2x^{2}-2$. For $q$ divisible by 33, it was shown in [26, Theorem 21] that ${R}^{\star}_{q}(1/2)\geqslant\log_{2}(q/3)$. A construction that is valid for all $q$ and that improves upon [26, Theorem 21] was given in [4, Section IV], showing that
${R}^{\star}_{q}(1/2)\;\geqslant\;\log_{2}\!\sqrt{\lfloor q/2\rfloor\!\cdot\!\lceil q/2\rceil}.$ (40)
In this section we provide a lower bound on ${R}^{\star}_{q}(1/2)$ that is valid for arbitrary $q$, and that coincides with (39) for $q=2$ and improves upon (40) for $q\in\{3,4\}$. It is based on a code construction that represents a generalization of the construction from [4, Section III] for the binary case.
Note that any lower bound on ${R}^{\star}_{q}(1/2)$ is automatically a lower bound on ${R}^{\star}_{q}(\tau)$ for all $\tau\in[0,\frac{1}{2}]$, and is potentially better than (30) for $\tau>\tau_{0}$ (see Figure 1). In other words, for any $\tau\in[0,\frac{1}{2}]$, we can write
${R}^{\star}_{q}(\tau)\;\geqslant\;\max\!\left\{R^{\textsc{gv}}_{q}(\tau),\,{R}^{\star}_{q}(1/2)\right\},$ (41)
where $R^{\textsc{gv}}_{q}(\tau)$ is the expression on the right-hand side of (30).
Construction:
Let $\{\mathcal{A}_{0},\mathcal{A}_{1}\}$ be a partition of the alphabet $\mathbb{Z}_{q}$, and
$\mathcal{P}\coloneqq\big\{(a,a,a),$ $(b,b,b),(a,b,b,b),(b,a,a,a),$ (42)
$(a,a,b,b,b,b),(b,b,a,a,a,a)\colon a\in\mathcal{A}_{0},b\in\mathcal{A}_{1}\big\}.$
Denote by $\mathcal{D}_{q}(n)$ the set of all strings of length $n$ that can be obtained as concatenations of the strings from $\mathcal{P}$.
Theorem 3.2
For every $q\geqslant 2$ and $n\geqslant 1$, we have $d_{\min}(\mathcal{D}_{q}(n))=\infty$, i.e., the code $\mathcal{D}_{q}(n)$ corrects an arbitrary number of transpositions. Consequently,
${R}^{\star}_{q}(1/2)\;\geqslant\;\log_{2}(\xi_{q}),$ (43)
where $\xi_{q}$ is the unique positive root of the polynomial $p_{q}(x)=x^{6}-qx^{3}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil x^{2}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$.
Proof
For the first part of the statement, we need to show that any codeword $\boldsymbol{x}\in\mathcal{D}_{q}(n)$ can be unambiguously recovered from the channel output $\boldsymbol{y}$ obtained after arbitrarily many transpositions have been applied to $\boldsymbol{x}$. Let $\boldsymbol{\tilde{x}}=(\tilde{x}_{1},\ldots,\tilde{x}_{n})$ denote the binary indicator string for $\boldsymbol{x}$ with respect to the partition $\{\mathcal{A}_{0},\mathcal{A}_{1}\}$, meaning that $\tilde{x}_{i}=j$ if and only if $x_{i}\in\mathcal{A}_{j}$, $j\in\{0,1\}$, and let $\widetilde{\mathcal{D}}(n)\coloneqq\{\boldsymbol{\tilde{x}}\colon\boldsymbol{x}\in\mathcal{D}_{q}(n)\}$. It was shown in [4, Theorem 3] that $\widetilde{\mathcal{D}}(n)$ is a binary code correcting an arbitrary number of transpositions. We then have the following situation: the receiver can infer $\boldsymbol{\tilde{y}}$ from $\boldsymbol{y}$ because it knows the sets $\mathcal{A}_{0},\mathcal{A}_{1}$, and it can infer $\boldsymbol{\tilde{x}}$ from $\boldsymbol{\tilde{y}}$ because $\boldsymbol{\tilde{x}}$ is a codeword of a transposition-correcting code. Hence, what is left for it to do is infer $\boldsymbol{x}$ from the pair $\boldsymbol{\tilde{x}}$, $\boldsymbol{y}$. It is easy to see, by analyzing the possible concatenations of blocks from (42), that this can always be done. First, by using $\boldsymbol{\tilde{x}}$ as the indicator string, all the transpositions in $\boldsymbol{y}$ that involve symbols from different sets, $a\in\mathcal{A}_{0}$, $b\in\mathcal{A}_{1}$, can be corrected. Second, the transpositions involving symbols from the same set, e.g., $a,a^{\prime}\in\mathcal{A}_{0}$, can also be easily recognized and corrected because they result in invalid blocks. For example, $(a,a,a,a^{\prime},a^{\prime},a^{\prime})$ may have resulted in $(a,a,a^{\prime},a,a^{\prime},a^{\prime})$ in the channel, $(a,a,a,a^{\prime},b,b,b)$ may have resulted in $(a,a,a^{\prime},a,b,b,b)$, etc. This means that the decoding process $\boldsymbol{y}\to(\boldsymbol{y},\boldsymbol{\tilde{y}})\to(\boldsymbol{y},\boldsymbol{\tilde{x}})\to\boldsymbol{x}$ is unambiguous, and hence the code $\mathcal{D}_{q}(n)$ corrects all patterns of transpositions.
It follows from the construction that the code cardinality satisfies the recurrence relation
$|\mathcal{D}_{q}(n)|=$ $|\mathcal{A}_{0}|\cdot|\mathcal{D}_{q}(n-3)|+|\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-3)|$ (44)
$+2|\mathcal{A}_{0}||\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-4)|+2|\mathcal{A}_{0}||\mathcal{A}_{1}|\cdot|\mathcal{D}_{q}(n-6)|.$
This implies that $|\mathcal{D}_{q}(n)|=\xi_{q}^{n+o(n)}$ as $n\to\infty$, where $\xi_{q}$ is the unique positive root of the characteristic polynomial $p_{q}(x)=x^{6}-(|\mathcal{A}_{0}|+|\mathcal{A}_{1}|)x^{3}-2|\mathcal{A}_{0}||\mathcal{A}_{1}|x^{2}-2|\mathcal{A}_{0}||\mathcal{A}_{1}|$. Taking $|\mathcal{A}_{0}|=\lfloor\frac{q}{2}\rfloor$, $|\mathcal{A}_{1}|=\lceil\frac{q}{2}\rceil$, the second part of the statement follows. $\blacksquare$
We note that the code rate obtained in Theorem 3.2 is larger than that in (40) for $q\leqslant 4$, but not for $q>4$.
Using (41), (40), and Theorems 3.1 and 3.2, we have, for any $q\geqslant 2$ and any $\tau\in[0,\frac{1}{2}]$,
${R}^{\star}_{q}(\tau)\;\geqslant\;\max\!\left\{R^{\textsc{gv}}_{q}(\tau),\,\log_{2}(\xi_{q}),\,\log_{2}\!\sqrt{\lfloor q/2\rfloor\!\cdot\!\lceil q/2\rceil}\right\},$ (45)
where $R^{\textsc{gv}}_{q}(\tau)$ is the expression on the right-hand side of (30), and $\xi_{q}$ is the unique positive root of the polynomial $p_{q}(x)=x^{6}-qx^{3}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil x^{2}-2\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$. The expression $\lfloor\frac{q}{2}\rfloor\lceil\frac{q}{2}\rceil$ simplifies to $\frac{q^{2}}{4}$ for even $q$, and $\frac{q^{2}-1}{4}$ for odd $q$.
The bound (45) is plotted in Figure 1 for the case $q=4$.
Figure 1: Lower bounds on the asymptotic rate of optimal codes correcting a fraction of $\tau$ transpositions over an alphabet of size $q=4$; see (45). The bound from Theorem 3.2 is better than that from Theorem 3.1 for $\tau>\tau_{0}\approx 0.172$.
4 Bounds on codes correcting $t=\textnormal{const}$ transpositions
In this section we present bounds on the cardinality of optimal codes in $\mathbb{Z}_{q}^{n}$ correcting $t=\textnormal{const}$ transpositions (Theorem 4.1 ahead). In contrast with the linear $t=\tau n$ regime, the lower bound obtained via the Gilbert–Varshamov argument is not better than the bounds that can be obtained by different methods, see Remarks 3 and 4 ahead. Nonetheless, we derive the expression for the average ball size and state the corresponding bound in the $t=\textnormal{const}$ regime as they may be of separate interest.
Lemma 1
For fixed $\mu\in[0,1]$,
$\lim_{n\to\infty}\frac{1}{n}\log_{q}\!\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\colon\operatorname{run}(\boldsymbol{x})\leqslant\mu n\}\right|=\begin{cases}H_{q}(\mu),&0\leqslant\mu<1-q^{-1}\\ 1,&1-q^{-1}\leqslant\mu\leqslant 1\end{cases},$ (46)
where $H_{q}(\cdot)$ is the $q$-ary entropy function.
Proof
Let $S_{q}(n,m)$ denote the number of strings from $\mathbb{Z}_{q}^{n}$ having exactly $m$ runs of identical symbols, so that
$\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\colon\operatorname{run}(\boldsymbol{x})\leqslant\mu n\}\right|=\sum_{m=1}^{\lfloor\mu n\rfloor}S_{q}(n,m).$ (47)
By a direct counting argument we find that
$S_{q}(n,m)=\binom{n-1}{m-1}q(q-1)^{m-1}$ (48)
and hence, by Stirling’s approximation,
$\lim_{n\to\infty}\frac{1}{n}\log_{q}S_{q}(n,\lfloor\mu n\rfloor)=H_{q}(\mu).$ (49)
Since the quantity we are interested in, expressed in (47), can be sandwiched between
$\max_{m\leqslant\lfloor\mu n\rfloor}S_{q}(n,m)\;\leqslant\;\sum_{m=1}^{\lfloor\mu n\rfloor}S_{q}(n,m)\;\leqslant\;\mu n\cdot\max_{m\leqslant\lfloor\mu n\rfloor}S_{q}(n,m),$ (50)
its exponential growth-rate (to base $q$) is
$\lim_{n\to\infty}\frac{1}{n}\log_{q}\sum_{m=1}^{\lfloor\mu n\rfloor}S_{q}(n,m)=\max_{0\leqslant\mu^{\prime}\leqslant\mu}H_{q}(\mu^{\prime}).$ (51)
That this is equal to the right-hand side of (46) follows from the fact that $H_{q}(\mu)$ is a concave function of $\mu\in[0,1]$ maximized at $\mu=1-q^{-1}$, the maximum being $H_{q}(1-q^{-1})=1$. $\blacksquare$
Lemma 1 implies that, for any fixed $\epsilon>0$, the number of strings having at most $(1-q^{-1}-\epsilon)n$ runs is exponentially smaller than the number of remaining strings, namely
$\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\colon\operatorname{run}(\boldsymbol{x})\leqslant(1-q^{-1}-\epsilon)n\}\right|=q^{nH_{q}(1-q^{-1}-\epsilon)+o(n)},$ (52)
where $H(1-q^{-1}-\epsilon)<1$. A very useful consequence of this fact is that one can disregard such strings in the asymptotic analysis of optimal codes in the $n\to\infty$, $t=\textnormal{const}$ regime. Namely, since $M^{\star}_{q}(n;t)$ is lower-bounded by the cardinality of optimal codes correcting $t$ substitutions (see Section 1.3), and the latter for constant $t$ scales as $\Omega(\frac{q^{n}}{n^{2t}})$ (by the standard Gilbert–Varshamov bound in the Hamming metric [15, Section 2.8]), we conclude that
$M^{\star}_{q}(n;t)=\Theta\!\left(q^{n+o(n)}\right).$ (53)
Relations (52) and (53) imply that, for an optimal code correcting $t$ transpositions, throwing out its codewords that have at most $(1-q^{-1}-\epsilon)n$ runs affects its cardinality by an asymptotically negligible amount.
Theorem 4.1
For any fixed $q\geqslant 2$ and $t\geqslant 1$, as $n\to\infty$,
$\frac{(2t)!\,q^{2t}}{(q-1)^{2t}}\frac{q^{n}}{n^{2t}}\;\lesssim\;M^{\star}_{q}(n;t)\;\lesssim\;\frac{t!\,q^{t}}{(q-1)^{t}}\frac{q^{n}}{n^{t}}.$ (54)
Proof
From Proposition 6 we can find the asymptotic behavior of $T^{\textnormal{e}}_{q}(n;r)$ for fixed $r$ and $n\to\infty$,
$T^{\textnormal{e}}_{q}(n;r)\;\sim\;n^{r}\frac{1}{r!}q^{n-r}(q-1)^{r},$ (55)
where we used the fact that $\binom{n}{s}\sim\frac{n^{s}}{s!}$ when $s$ is fixed and $n\to\infty$. From (55), (17) and (18) it follows that $T_{q}(n;r)$ and $\bar{T}_{q}(n;r)$ have the same asymptotic scaling as $T^{\textnormal{e}}_{q}(n;r)$ in the regime currently of interest. This fact, together with the bound ${M}_{q}^{\star}(n;t)\geqslant\frac{q^{2n}}{\bar{T}_{q}(n;2t)}$ (see (31)) implies the left-hand side of (54).
In order to show the right-hand inequality, an asymptotic lower bound on the cardinality of $B(\boldsymbol{x};t)$ (the set of all strings that can be obtained by applying at most $t$ transpositions to the string $\boldsymbol{x}$) will be needed. Using Proposition 4, for strings satisfying $\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n$, $|B(\boldsymbol{x};t)|$ can be asymptotically lower-bounded by
$\gtrsim\;\sum_{u=0}^{t}\binom{(1-q^{-1}-\epsilon)n/2}{u}\binom{(1-q^{-1}-\epsilon)n/2}{t-u}$ (56)
$\sim\;\sum_{u=0}^{t}\frac{(1-q^{-1}-\epsilon)^{u}\,n^{u}}{2^{u}\,u!}\frac{(1-q^{-1}-\epsilon)^{t-u}\,n^{t-u}}{2^{t-u}\,(t-u)!}$
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{2^{t}}\,\sum_{u=0}^{t}\frac{1}{u!\;(t-u)!}$
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{2^{t}\,t!}\,\sum_{u=0}^{t}\binom{t}{u}$
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{t!}.$
We can now derive the upper bound in (54) by a packing argument. Let $\mathcal{C}^{\star}_{q}(n;t)$ be an optimal code correcting $t$ transpositions, meaning that $|\mathcal{C}^{\star}_{q}(n;t)|=M^{\star}_{q}(n;t)$. By (3), the sets $B(\boldsymbol{x};t)$ centered at codewords of $\mathcal{C}^{\star}_{q}(n;t)$ do not overlap, so we have
$\sum_{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n;t)\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n}|B(\boldsymbol{x};t)|\;\leqslant\;\sum_{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n;t)}|B(\boldsymbol{x};t)|\;\leqslant\;|\mathbb{Z}_{q}^{n}|\;=\;q^{n}.$ (57)
Together with (56) this implies that
$\frac{1}{t!}\left(1-q^{-1}-\epsilon\right)^{\!t}n^{t}\cdot\left|\left\{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n;t)\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n\right\}\right|\;\lesssim\;q^{n},$ (58)
From Lemma 1 and the discussion thereafter we know that we can safely ignore in the asymptotic analysis the codewords with at most $(1-q^{-1}-\epsilon)n$ runs, meaning that
$\left|\left\{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n;t)\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n\right\}\right|\;\sim\;\left|\mathcal{C}^{\star}_{q}(n;t)\right|\;=\;M^{\star}_{q}(n;t).$ (59)
Now (58) and (59) imply
$\frac{1}{t!}\left(1-q^{-1}-\epsilon\right)^{\!t}n^{t}\cdot M^{\star}_{q}(n;t)\;\lesssim\;q^{n}.$ (60)
Since (60) holds for any $\epsilon>0$, the upper bound in (54) is established. $\blacksquare$
Remark 3(Lower bound for the binary case)
For $q=2$, a lower bound better than the one in Theorem 4.1 was derived in [21, Theorem 11] by using an interpretation of binary codes correcting successive transposition errors as packings in the Manhattan metric, and using the densest known packings, namely
$M^{\star}_{2}(n;t)\;\gtrsim\;c_{t}\frac{2^{n+t}}{n^{t}},$ (61)
where $c_{1}=\frac{1}{2}$, $c_{2}=\frac{1}{3}$,333[21, Theorem 11] states a slightly worse constant $c_{2}=\frac{1}{4}$. It can be shown based on [30, Theorem 3] that $c_{2}$ can be improved to $\frac{1}{3}$. and $c_{t}=\frac{1}{2t+1}$ for $t\geqslant 3$. In particular, in the binary case we have
$M^{\star}_{2}(n;t)=\Theta\Big(\frac{2^{n}}{n^{t}}\Big).$ (62)
Unfortunately, as already mentioned in Remark 2, this geometric interpretation fails when $q>2$ and thus the construction that implies this lower bound does not generalize to larger alphabets. $\blacklozenge$
Remark 4(Lower bound via substitution-correcting codes)
For general $q$, a lower bound better than the one in (54) follows from the known bounds for substitution-correcting codes (see Section 1.3). For example, it is known that $q$-ary codes correcting a single substitution and having cardinality $\Theta(\frac{q^{n}}{n})$ exist [15] (e.g., Hamming codes), and hence
$M^{\star}_{q}(n;1)\;=\;\Theta\Big(\frac{q^{n}}{n}\Big).$ (63)
For $t>1$, Varshamov’s improvement of Gilbert’s bound [15, Section 2.9] implies that
$M^{\star}_{q}(n;t)\;=\;\Omega\Big(\frac{q^{n}}{n^{2t-1}}\Big).$ (64)
Further improvements of the exponent in the denominator are possible in some cases; see, e.g., [8, 34].
Based on what is known in the special case $t=1$ (see (63)), as well as in the case of general $t$ and $q=2$ (see (62)), it is reasonable to conjecture that
$M^{\star}_{q}(n;t)\;\stackrel{{\scriptstyle?}}{{=}}\;\Theta\Big(\frac{q^{n}}{n^{t}}\Big),$ (65)
i.e., that the redundancy of $t\log_{q}n+{\mathcal{O}}(1)$ is achievable for any fixed $t$ and alphabet size $q$. At present, however, we do not have a proof of this fact. $\blacklozenge$
5 Conclusion and further work
We have analyzed the problem of correcting transpositions of consecutive symbols in $q$-ary strings—types of errors that may occur in various contexts, e.g., as synchronization and timing errors, spelling errors, inter-symbol interference errors, genome rearrangement errors, etc.
In the linear regime $t=\tau n$, we have presented a Gilbert–Varshamov-like lower bound on the achievable code rates. Since the code space is nonuniform, obtaining good upper bounds is a nontrivial problem that is left for future work, as are explicit constructions of asymptotically good (i.e., having positive rate) codes correcting a fraction of $\tau$ transpositions. For codes of infinite minimum distance, the problem of deriving the exact value of the maximum achievable rate, i.e., the zero-error capacity, is very interesting and is still open for all alphabet sizes. In the $t=\textnormal{const}$ regime, as one of the main open questions in this line of work we mention a construction of nonbinary codes correcting $t>1$ transpositions whose cardinality scales as $\Theta(\frac{q^{n}}{n^{t}})$ when $n\to\infty$ (see (65)).
Finally, there are several related models describing communication in the presence of synchronization and reordering errors that are potentially relevant for modern information transmission and storage systems and that are not yet fully understood, e.g., timing channels, permutation channels with bounded displacements, channels with a combination of transpositions and other types of errors, etc.
5.0.1 Acknowledgments.
M. Kovačević was supported by the Ministry of Science, Technological Development and Innovation of the Republic of Serbia (contract no. 451-03-137/2025-03/200156) and by the Faculty of Technical Sciences, University of Novi Sad, Serbia (project no. 01-50/295).